### v2

# Transcendence of ${}_{3}F_{2}(1)$ Hypergeometric Series and L-values of Modular Forms

Abstract.
We examine a certain class of hypergeometric motives, recently associated to Hecke eigenforms of weight 3 by Allen, Grove, Long, and Tu. One aspect of their results allows us to express a special $L$-value of a specific Hecke eigenform as a combination of ${}_{3}F_{2}(1)$ hypergeometric series. First, we show that the $L$-function can be twisted in a predictable way so that the $L$-value of the twist can be also be written as a sum of hypergeometric series. Moreover, the hypergeometric series involved in the twist are related to a classical formula of ${}_{3}F_{2}(1)$ series due to Kummer, which we use the extend to relations between $L$-values of the eigenforms. In the process, we find a bound on the transcendence degree for certain ${}_{3}F_{2}(1)$ series. This suggests a generalization of a theorem due to Wüstholz which provides a formula for the transcendence degree of periods on abelian varieties.
1. Introduction
There is a long history of studying hypergeometric series and their transformation formulas, going back the work of Gauss, Kummer, and many others. For instance, the central transformation formula to this paper is the following.
Theorem 1.1 (Kummer, 1836).
When both series converge
[…]
The version here is a special case of the more general theorem; see Andrews, Askey, and Roy [2] Cor. 3.3.5. Since Greene [19] introduced finite field hypergeometric functions, a great deal of work has focused on extending these identities to the finite field setting. However, there are numerous other ways the identities can be interpreted as well, especially when modular forms are in the background. Some motivating examples are [11], [24], and [42].
Recently, Allen, Grove, Long, and Tu [1] studied the relation between the hypergeometric motive attached to
$HD(r,s):=\{\{1/2,1/2,r\},\{1,1,s\},\{1\}\}$
for certain choices of rational $r,s$, and modular forms. Among other results, they showed that a normalized—see (8)— classical hypergeometric series ${}_{3}F_{2}(HD(r,s))$ is the exact $L$-value at 1 of an explicit weight 3 modular form denoted $\mathbb{K}_{2}(r,s)$ to be defined below (5). For instance,
[…]
We observe that in Kummer’s formula above, that then the $L$-value in terms of ${}_{3}F_{2}(HD(r,s))$ transforms to a multiple of ${}_{3}F_{2}(HD(r^{\prime},s^{\prime}))={}_{3}F_{2}(HD(1-r,1/2-r+s))$. This implies a relation between the $L$-values of the form
(1) $L(\mathbb{K}_{2}(r,s),1)=\alpha_{r,s}\cdot L(\mathbb{K}_{2}(r^{\prime},s^{\prime}),1).$
Generally, $\mathbb{K}_{2}(r,s)$ is not a Hecke eigenform. However, in [1], they discuss several examples where $\mathbb{K}_{2}(r,s)$ can be used to construct a Hecke eigenform, $f_{r,s}$, unique up to a quadratic twist, which we fix in Table 6. In these special cases, which we call Galois (Definition 2.3), we can use the action of the Hecke operators to construct $f_{r,s}$ using $\mathbb{K}_{2}(r,s)$ (see Theorem 3.4). The construction relies on the notion of conjugates given in [1] (Definition 2.2), motivated by the idea of Galois conjugates. For instance, if we start with the function $\mathbb{K}_{2}(1/8,5/8)$, the conjugates are $\mathbb{K}_{2}(i/8,i/8+1/2)$ for $i=3,5,7$. If $M$ is the lowest common denominator of $1/2,r,$ and $s$, the cyclotomic field is $\mathbb{Q}(\zeta_{M})$, where $\zeta_{M}$ is a primitive $M$-th root of unity, play a crucial role in the arithmetic of the hypergeometric function. From high level point of view, this is because the hypergeometric motive associated to $HD(r,s)$ is defined over $\mathbb{Q}(\zeta_{M})$. In our example, this is reflected in that each of the three conjugates corresponds to a Galois conjugate of $\zeta_{8}$ over $\mathbb{Q}$. Refer to Appendix II for more details on the geometric situation.
[…]
In the Part 2, we prove the extended version of the Kummer transformation is
(2) $L(f_{1/8,5/8},1)=-\frac{1}{2}\zeta_{48}\cdot L(f_{1/8,1},1).$
The explanation of this theorem is given in terms of modular forms, which is stated using the Fourier expansion of the modular form at $i\infty$. We write $q=e^{2\pi i\tau}$. Then, there is a striking relationship between the conjugate families $\{(i/4,1)\text{ }|\text{ }i=1,3\}$ and $\{(i/4,i/4+1/2),\text{ }|\text{ }i=1,3\}$: the Fourier coefficients are
$\mathbb{K}_{2}(1/4,1)=q-2q^{5}-7q^{9}+14q^{13}+18q^{17}-32q^{21}-21q^{25}+14q^{29}+O(q^{30})$
$\mathbb{K}_{2}(1/4,3/4)=q+2q^{5}-7q^{9}-14q^{13}+18q^{17}+32q^{21}-21q^{25}-14q^{29}+O(q^{30}),$
that is, they differ by a sign. A similar property holds for $\mathbb{K}_{2}(3/4,1)$ and $\mathbb{K}_{2}(3/4,5/4)$. Moreover, the Kummer transformation interchanges the $L$-values of these families of forms as in (1). This general twisting property is our first main theorem, stated more precisely in Theorem 3.6.
Theorem 1.2.
Let $h=h(r,s)=r-s+3/2$. If $20<h(r,s)<3/$ 2, then for any positive integer $n$, the $n$th Fourier coefficients of $\mathbb{K}_{2}(r,s)(\tau)$ and $\mathbb{K}_{2}(r,h)(\tau)$ are the same up to a sign. Moreover, if $(r,s)$ belongs to a Galois family, then the completed Hecke eigenforms $f_{r,s}$ and $f_{r,h}$ differ by a finite order character.
In our example above, we have have $f_{1/4,1}=f_{1/4,3/4}\otimes\chi_{1/4,3/4}$ for some finite order character. Likewise, $f_{1/8,1}=f_{1/8,5/8}\otimes\chi_{1/8,5/8}$ In our examples above, $\chi$ is an even character, and so a theorem of Shimura [30], which we state as Theorem 4.5 below, explains (2). Our example above is one of the examples in Theorem 4.6 and is rewritten as
(3) $L(f_{1/8,5/8},1)=-\frac{1}{2}\zeta_{48}\cdot L(f_{1/8,5/8}\otimes\chi_{1/8,5/8},1).$
Shimura’s theorem predicts relations such as these, completing the story of the Kummer transformation’s relation with $L$-values in our setting.
[…]
There is a long history of studying the transcendence of important numbers such as $\pi$ and $e$. Some well-known examples are studied in [9]. However, transcendence problems are generally very hard. For example, the values of the Riemann zeta function at odd integers are anticipated to be transcendental, yet so far, only the irrationality of $\zeta(3)$ is definitively known due to a famous result of Apéry, which has also been interpreted in terms of modular forms by Beukers [6]. Some progress has been made by Zudilin and his collaborators; for example, in [41] he shows that one of $\zeta(5),\zeta(7)$, $\zeta(9)$, and $\zeta(11)$ is irrational. However, proving irrationality for a given odd zeta value other than 3 remains open. Interestingly, irrationality results about the zeta function can be related to certain integrals of hypergeometric type, see [41] as well as [25] by Rhin and Viola.
The motivating result applicable to our situation is a fundamental theorem of Wüstholz [36], a corollary of the Analytic Subgroup Theorem, which is stated in terms of general algebraic groups — see [29] for a formulation in terms of abelian varieties. We state a specific case relevant to our situation.
Theorem 1.3 (Wüstholz).
The set of periods of differentials of the first kind on a simple abelian surface $A$ with quaternionic multiplication (QM) over $\overline{\mathbb{Q}}$ is 2.
Recall an abelian variety has quaternionic multiplication if the endomorphism algebra of the variety contains a quaternion algebra. In other words, if an abelian variety has extra endomorphisms, then this extra structure reduces the dimension of the vector space above. Wüstholz’s theorem is central to many other transcendence results, some of which are discussed in the survey paper [34] by Waldschmidt.
An application of our three-term relations suggests a higher dimensional version of Wüstholz’s theorem. We replicate the QM situation above in the following way. Let $M$ be the least common denominator of $1/2,r,s$ as above. By work of Katz [21], there is a 2-dimensional Galois representation $\sigma_{HD(r,s)}$ associated to $HD(r,s)$ defined over the absolute Galois group $G_{\mathbb{Q}(\zeta_{M})}$, unramified at all but finitely many primes of $\mathbb{Q}(\zeta_{M})$, whose trace of Frobenius at unramified primes is related to a finite field analogue of hypergeometric functions. Throughout, we use $\varphi(M)$ to denote the Euler phi function evaluated at $M$. Suppose for some number field $K$ the induction
(4) $\rho_{HD}:=\text{Ind}_{G_{\mathbb{Q}(\zeta_{M})}}^{G_{\mathbb{Q}}}\sigma_{HD}$
satisfies
$\rho|_{G_{K}}=\sigma_{1}\oplus\sigma_{2}\oplus\dots\oplus\sigma_{\varphi(M)},$
where $\rho|_{G_{K}}$ denotes the restriction to the absolute Galois group $G_{K}$ and the $\sigma_{i}$ are 2-dimensional and isomorphic to each other. In this case, we say $\sigma_{HD}$ is potentially 2-isotypic (see [23]). Using ${}_{2}F_{1}$ series in [14], they observed that QM abelian surfaces also have 2-dimensional Galois representations which are potentially 2-isotypic, so we may view this as a generalization of the 2-dimensional abelian varieties admitting QM. There is also a complex analytic interpretation given in Appendix II.
Our transcendence result is the following analogue of Wüstholz’s theorem.
Theorem 1.4.
Assume Conjecture 4.11 below, and that $M=\text{lcd}(r,s,1/2)$ as above. For the non-CM Galois families satisfying $\varphi(M)\leq 4$, the vector space for periods of differentials of the first kind over $\bar{\mathbb{Q}}$ has dimension less than or equal to $2$.
Conjecture 4.11 is a mild transcendence results that is required for two out of the eight cases addressed by Theorem 1.4.
The paper is organized in two main parts. Assuming Corollary 3.3 and Theorem 3.6, the two parts can be read independently. First, we discuss the necessary background in Section 2. Part 1 then focuses on properties of the $\mathbb{K}_{2}(r,s)(\tau)$, classifying all the possible families (Table 1 and Table 2), and setting up the necessary tools to study the $L$-functions of Hecke eigenforms, $f_{r,s}$. The first half of this Part is dedicated to showing the action of the Hecke operators on Galois families in order to methodically construct $f_{r,s}$. Then, we prove Theorem 1.2 and and use that to classify all possible Galois families.
In Part 2, we prove the extension of the Kummer transformation to the $L$-values of Hecke eigenforms. In the process, we prove Theorem 1.4. After discussing some further directions, we provide a table of all Hecke eigenforms obtained using our method in Appendix I. In Appendix II, we provide some geometric explanation for the transcendence result.
Throughout, we focus exclusively on the situation where $f_{r,s}$ is non-CM. More details about CM examples, as well as non-Galois examples, will be provided in a forthcoming paper, [27].
Acknowledgments
The author would like to express her gratitude to Ling Long and Fang-Ting Tu for suggesting this problem based on their recent work, and for their helpful guidance and suggestions along the way. She would also like to thank John Voight and Wadim Zudilin for helpful discussions. Finally, she thanks Michael Allen, Brian Grove, and Hasan Saad for their helpful comments on an earlier version of this paper.
The author is supported by a summer research assistantship from the Louisiana State University Department of Mathematics.
2. Preliminaries
In this section, we introduce the $\mathbb{K}_{2}(r,s)$ functions and recall some of their important properties from [1]. We also briefly explain how we can frame our discussion throughout the paper using a finite Coxeter group.
2.1. The $\mathbb{K}_{2}(r,s)$ Functions
For rational parameters $a_{i}$ and $b_{i}$ and $\lambda\in\mathbb{C}\setminus\{0,1\}$, let $(a)_{k}=a(a+1)\dots(a+k-1)$. Then define the generalized hypergeometric series by
$\,{}_{n}F_{n-1}\left[\begin{matrix}a_{1}&a_{2}&\dots&a_{n}\vskip 3.0pt plus 1.0pt minus 1.0pt\\ b_{1}&b_{2}&\dots&b_{n-1}\end{matrix}\;;\;\lambda\right]=\sum_{k=0}^{\infty}\frac{(a_{1})_{k}(a_{2})_{k}\dots(a_{n})_{k}}{(b_{1})_{k}(b_{2})_{k}\dots(b_{n-1})_{k}}\frac{\lambda^{k}}{k!}.$
It is well known (see e.g. [37]) that when $\lambda$ is the modular lambda function, the hypergeometric function
$\,{}_{2}F_{1}\left[\begin{matrix}1/2&1/2\vskip 3.0pt plus 1.0pt minus 1.0pt\\ &1\end{matrix}\;;\;\lambda\right]$
is a weight 1 modular form on the congruence group $\Gamma(2)$. Let $q=e^{2\pi i\tau}$. Multiplying by $d/dq\log\lambda=d\lambda/(2\pi i\lambda)$ then yields a weight 3 weakly holomorphic modular form, which can be made holomorphic by multiplying appropriate powers of $\lambda$ and $1-\lambda$. Using this idea, Allen, Grove, Long, and Tu [1] defined the function
(5) $\mathbb{K}_{2}(r,s)(\tau)=\frac{\eta(\tau/2)^{16s-8r-12}\eta(2\tau)^{8r+8s-12}}{\eta(\tau)^{24s-30}}.$
They also showed there are finitely many $(r,s)$ which correspond to holomorphic congruence modular forms, and the set of all of these pairs is
$\mathbb{S}_{2}^{\prime}:=\{(r,s)\text{ }|\text{ }0<r<s<3/2,\text{ }24s,8(r+s)\in\mathbb{Z}\}.$
[…]
The $\mathbb{K}_{2}(r,s)$ functions are closely related to hypergeometric functions by the following Lemma.
Lemma 2.1 ([1]).
The function $\mathbb{K}_{2}(r,s)$ is a weight 3 cusp form for some finite index subgroup $\Gamma$ of $SL_{2}(\mathbb{Z})$ if and only if $20<r<s<3/$ 2, and
$\mathbb{K}_{2}(r,s)(\tau)=2^{1-4r}\lambda^{r}(1-\lambda)^{s-r-1}\,_{2}F_{1}\left[\begin{matrix}1/2&1/2\vskip 3.0pt plus 1.0pt minus 1.0pt\\ &1\end{matrix}\;;\;\lambda\right]\frac{d}{dq}\log\lambda.$
Moreover, $\Gamma$ is congruence if and only if the exponents in the eta product (5) are integral, and the level is
$\frac{48}{\gcd(24r,24)}\frac{48}{\gcd(24(s-r),24)}.$
For the remainder of this paper, we will only work with $\mathbb{K}_{2}(r,s)$ defined on congruence subgroups. It is well established that such modular forms have a Fourier expansion in terms of the variable $q$. The $q$-expansion of $\mathbb{K}_{2}(r,s)(\tau)$ should a priori have local uniformizer $e^{2\pi i\tau/N}$, where
(6) $N=\frac{48}{\gcd(24r,24)}.$
However, it turns out we can lift all of our congruence $\mathbb{K}_{2}(r,s)(\tau)$ to $\Gamma_{1}(\mathcal{N})$ by using $\mathbb{K}_{2}(r,s)(N\tau)$, for which the local uniformizer is $q=e^{2\pi i\tau}$ as above. Despite this, $N$ appears in several formulas later. Note that by construction, if $r=a/b$ where $a$ and $b$ are coprime, then $N=2b$. We will use this fact several times. By a slight abuse of notation, when we write $\mathbb{K}_{2}(r,s)$ without a variable, we mean $\mathbb{K}_{2}(r,s)(N\tau)$.
The modular forms $\mathbb{K}_{2}(r,s)$ are also related to generalized hypergeometric series via their $L$-values. More precisely, the integral representation of hypergeometric functions (see [2]) gives that
[…]
We also have the integral representation for the special $L$-values attached to $\mathbb{K}_{2}(r,s)$,
[…]
For convenience, we will establish the notation
[…]
Remark 1.
Our normalization is non-standard and is chosen so that the computations with $L$-values are more convenient. Note that it differs from the $P(r,s)$ function used in [1] as follows
$\pi F(r,s)=N2^{1-4r}P(r,s).$
The only notable difference is the multiple of $\pi$ from a transcendence perspective. The factor of $\pi$ does not impact our main results, which are stated as linear relations, where each term in the relation should be multiplied by $\pi$ to get $P(r,s)$.
Note, $\pi F(r,s)$ is a period on a hypergeometric surface, which is important from a geometric point of view. If $F(r,s)$ and $F(r^{\prime},s^{\prime})$ are conjugate, then they can be defined as periods on the same surfaces - see Lemma 7.1 in Appendix II. This reduces to the following definition. As above, $M=\text{lcd}(1/2,r,s)$, where we assume $r$ and $s$ are simplified into lowest terms.
Definition 2.2.
Two pairs $(r,s)$ and $(r^{\prime},s^{\prime})$ are conjugate if there exists an integer $c$ coprime to $M$ so that $r-cr^{\prime}$ and $s-cs^{\prime}$ are both integers.
While the conjugates put the $\mathbb{K}_{2}(r,s)$ functions into families based on the underlying hypergeometric varieties, we can also use information about them in modular forms.
Definition 2.3.
A family of conjugate pairs $(r_{i},s_{i})$ is Hecke stable if each $\mathbb{K}_{2}(r_{i},s_{i})$ is in the same Hecke orbit for all $i$. If for $i\neq j$, $r_{i}\neq r_{j}$, then we call a the family Galois
In the Galois case, we can write our family as
$S_{i/b,s_{i}}:=\{\mathbb{K}_{2}(i/b,s_{i})\}_{i\in(\mathbb{Z}/b\mathbb{Z})^{\times}}.$
We will use this extensively in Section 3.1.
The Galois condition is crucial since it ensures that the induction $\rho$ of $\sigma_{HD}$ given in (4) is potentially 2-isotpyic, as in (4). In particular, this means the finite field versions of the periods are each others Galois conjugates, allowing the $G_{\mathbb{Q}(\zeta_{M})}$-representation to be extended to $G_{\mathbb{Q}}$ (see[1]). This is reflected in the additional symmetries among the periods, which will be our focus later. We can show a family is Hecke stable by computing the action of the Hecke operators.
(10) $T_{p}\mathbb{K}_{2}(1/8,n/8+1/2)=C_{p}\mathbb{K}_{2}(p/8,p/8+1/2)$
for some integer $C_{p}$, so these conjugates are in the same Hecke orbit. Clearly, the family is also Galois. A similar property holds for all Galois families (see Lemma 3.2 below). Interestingly, there are cases which are not Hecke stable as well. These will be addressed in detail in the forthcoming paper [27]. Identifying the Hecke orbit of the modular forms in the LMFDB is possible since we know the level from Lemma 2.1, and so this is a finite check.
2.2. Coxeter Group Interpretation
There is a long history of studying hypergeometric transformations in terms of symmetries of a finite Coxeter group going back to at least Bailey [5]. A finite Coxeter group is really just a special kind of finite group, which for us will typically be $S_{n}$, the symmetric group on $n$ variables, or $D_{n}$, the dihedral group of order $2n$. For ${}_{3}F_{2}(1)$ series, this has been studied in great depth by Beyer et al. [8], and there are similar results for certain ${}_{4}F_{3}(1)$ [17], [18] as well. Using the finite symmetry group is very helpful for our classification of $\mathbb{K}_{2}(r,s)$ families and for stating other identities. This is because the Coxeter group provides a uniform way to state identities that come from the hypergeometric side (such as analytic continuation formulas) and the modular forms side (such as the Atkin–Lehner involution).
Based on the work of Bailey, Beyer et al. show that for ${}_{3}F_{2}(1)$ series, the invariance group for a properly normalized series denoted $\hat{F}(r,s)$ is $S_{5}$, and they use this to classify all two term identities between the hypergeometric series, such as the Kummer transformation. The action of $S_{5}$ on a hypergeometric multiset $\{\{a,b,c\},\{d,e\}\}$ is expressed by writing the multiset as a column vector $(a,b,c,d,e)^{t}$. Denote such a vector by $\mathfrak{a}$. Then we write $S_{5}$ as an appropriate subgroup of $GL_{5}(\mathbb{Z})$ so that our action on $\mathfrak{a}$ is the usual multiplication of a matrix and a vector. From the point of view of the Coxeter groups, the initial multiset $HD$ is not so important, but we will assume that $a=b=1/2,d=1$ for convenience throughout. Thus, we may write $\mathfrak{a}=(r,s)$ in our setting, with $a,b$ and $d$ implicit, so $F(r,s)=F(\mathfrak{a})$. The Kummer transformation 1.1 in our normalization is
$F(\mathfrak{a})=\alpha_{r,s}\cdot F(K\mathfrak{a})$
for $K$ the $5\times 5$ matrix in the proof of Lemma 2.4 and $\alpha_{r,s}$ a gamma quotient given explicitly in (17) below.
[…]
${}_{3}F_{2}(\mathfrak{a})(1)=\alpha\cdot{}_{3}F_{2}(m_{1}\mathfrak{a})(1)+\alpha^{\prime}\cdot{}_{3}F_{2}(m_{2}\mathfrak{a})(1)$
where $\alpha,\alpha^{\prime}$ are constants and $m_{1}$ and $m_{2}$ are certain $5\times 5$ matrices specified in [8]. A more explicit realization of this result is given in Proposition 4.12 below. They also show that for three-term identities, the invariance group is $S_{6}\times C_{2}$, where $C_{2}$ is the group of order 2, but the process of producing identities is more complicated. We will discuss three-term identities extensively in the second half of the paper.
There are two natural actions on the modular forms $\mathbb{K}_{2}(r,s)$, the Kummer transformation (Theorem 1.1) discussed above and the Atkin–Lehner involution. We will only use $W_{2}$, which is defined by the map $W_{2}:\tau\mapsto-1/2\tau$. Recall if $r=a/b$ is reduced, then $N=2b$. The action of $W_{2}$ is
$W_{2}\mathbb{K}_{2}(r,s)(2b\tau)=\tau^{-3}\text{ }\mathbb{K}_{2}(r,s)(-b/\tau)=2^{8s-16r}(-i)^{3}\cdot\mathbb{K}_{2}(s-r,s)(b\tau).$
This is straightforward to prove using the well-known formula $\eta(-1/\tau)=\sqrt{-i\tau}\eta(\tau)$. We fix the identity embedding of $\sqrt{-1}$ into $\mathbb{C}$, which we denote by $i$, for the remainder of this paper. The action of $W_{2}$ and the Kummer transformation on $F(r,s)$ are given by the operators
(11) $A:(r,s)\mapsto(s-r,s)$
and
(12) $K:(r,s)\mapsto(1-r,1/2-r+s).$
Proposition 2.4.
The group $G$ generated by $A$ and $K$ is isomorphic to $D_{6}$, the dihedral group of order 12.
Proof.
Both $A$ and $K$ are involutions, i.e., they have order two.
The general version of Theorem 1.1 given in Corollary 3.3.5 of [2] gives
$K((a,b,c,d,e)^{t})=(a,d-b,d-c,d,d+e-b-c)^{t}.$
Thus, with respect to the standard basis for $\mathbb{Q}^{5}$, the matrix for $K$ is
$K=\begin{pmatrix}1&0&0&0&0\\ 0&-1&0&1&0\\ 0&0&-1&1&0\\ 0&0&0&1&0\\ 0&-1&-1&1&1\end{pmatrix}.$
The matrix for $A$ is defined similarly. We also compute that
$AK=\begin{pmatrix}1&0&0&0&0\\ 0&-1&0&1&0\\ 0&-1&0&0&1\\ 0&0&0&1&0\\ 0&-1&-1&1&a\end{pmatrix}.$
Using linear algebra, it is then easy to check that $A$ and $K$ have order 2 and $AK$ has order 6, and so the group is isomorphic to the dihedral group of order 12. ∎
Note $D_{6}$ is a subgroup of the group $S_{6}\times C_{2}$ given in [8] for three term identities. We also point out that $D_{6}$ as a group is a subgroup of $S_{5}$ as well; however, using the normalization of [8], denoted $\hat{F}(r,s)$, we should not view our group $G$ as a subgroup of the invariance group $S_{5}$. Note that $\hat{F}(K(r,s))=\hat{F}(r,s)$ from [8], and so $K$ is in the invariance group of $\hat{F}(r,s)$. However, $\hat{F}(A(r,s))\neq\hat{F}(r,s)$ in general. For example, we can see $A(1/8,5/8)=(1/2,5/8)$, and it is easy to check numerically that $\hat{F}(1/8,5/8)\neq\hat{F}(1/2,5/8)$. This implies $A$ is not an element of the invariance group. As a result, besides the matrix $K$ arising from the Kummer transformation, we will use the group $G$ exclusively for three-term identities.
3. Part 1: Symmetries from Modular Forms and Coxeter Groups
In this section, we first show the technical but crucial fact that every Galois family of conjugates can be completed to a Hecke eigenform. This allows us to compute exact $L$-values of the modular forms and construct Table 3 in the Appendix. Next, we prove Theorem 1.2 and provide a classification of all families.
3.1. Building Eigenforms from the $\mathbb{K}_{2}$ Functions
3.1.1. Hecke Operators
First, we prove that the Hecke operators can be used to construct Hecke eigenforms out of Galois families. Denote the newform for the Hecke orbit of our Galois family by $f_{r,s}$, its character by $\psi_{r,s}$, and its Hecke eigenvalue field by $K_{f}$.
In Table 1, we organize families using an element of the form $(1/b,s^{\prime}/m)$,and we specify the quadratic twist $f_{1/b,s}$ on Table 3. Thus, we prove that such a pair always exists exists. As above, and throughout, we will assume that $a$ and $b$ are coprime.
Lemma 3.1.
Assume $b\geq 2$ is an integer and $a$ is an integer coprime to $b$. Then every $\mathbb{K}_{2}(a/b,s/m)$ defined on a congruence subgroup for $a,b,$ and $s$ integers is conjugate to $\mathbb{K}_{2}(1/b,s^{\prime}/m)$ for some integer $21/b<s^{\prime}/m<3/$ 2.
Proof.
The case $b=2$ is trivial, since the only possible $a/2$ lying between 00 and $3/2$ is 1/2. So assume $b>2$. Note $(1/b)(b-a)+a/b$ equals 1, an integer. Also observe that $b-a$ is coprime to $M$, since $(a,b)=1$. So we only need to prove there is some $s^{\prime}$ so that $(b-a)s^{\prime}/m+s/m$ is also an integer, that is, $(b-a)s^{\prime}+s\equiv 0\mod m$. For the pair $(a/b,s/m)$ to correspond to a modular form on a congruence group, $m$ has to be either 1, 2, 4, 6, 8, 12, or 24, all of which can be written as products consisting of powers of 2 and 3. Therefore, $(b-a,m)=1$, and it follows that the linear congruence
(13) $(b-a)s^{\prime}+s\equiv 0\mod m$
has a unique solution modulo $m$. We choose a solution $x$ of (13) so that $x\leq m$. If $x/m\geq 1/b$ we only need $x/m<3/2$ so that the conjugate $(1/b,x/m)$ constructed above is holomorphic. Because we assumed $x\leq m$, we have $x/m\leq 1<3/2$. Therefore, we can take $s^{\prime}=x$. When $x/m<1/b$, since $b>2$, we have $x/m<1/2$, and it follows that $21/n<x/m+1<3/$ 2. Moreover, clearly $x+m$ satisfies the congruence (13) as well, so we take $s^{\prime}=x+m$. ∎
Remark 2.
Each $\mathbb{K}_{2}(a_{i}/b,s_{i})$ has nonzero $n$th Fourier coefficients only if $n\equiv a_{i}\mod b$. This can be checked by directly expanding the eta product definition of $\mathbb{K}_{2}(a_{i}/b,s_{i})(N\tau)$.
Lemma 3.2.
Let $\{\mathbb{K}_{2}(i/b,s_{i})\}_{i\in(\mathbb{Z}/b\mathbb{Z})^{\times}}$ be a Galois family. Let $p$ be a prime so that $p\equiv i\mod b$. Then there exists a nonzero integer $C(p,j)$ so that for any $j\in(\mathbb{Z}/b\mathbb{Z})^{\times}$,
$T_{p}\mathbb{K}_{2}(j/b,s_{j})=C(p,j)\cdot\mathbb{K}_{2}(k/b,s_{k}),$
where $ij=k$ as elements of $(\mathbb{Z}/b\mathbb{Z})^{\times}$. In other words, the action of the Hecke operators is governed by the group structure of $(\mathbb{Z}/b\mathbb{Z})^{\times}$.
Remark 3.
Lemma 3.2 is a more general version of an example given in Corollary 3.1 of [1] for the family $(1/8,1)$. They can use the Dwork dash operator since $s=1$, which has the same action. In this sense, Lemma 3.2 tells us that we can use the Hecke operators to generalize the Dwork dash operator, which does not apply when $s\neq 1$, to Galois families.
Proof.
Note that $S_{i/b,s_{i}}$ generates a vector space with dimension larger than the dimension of $K_{f}$. This is because the set spans the Hecke orbit and the $q$-coefficients are all integers. In particular, the set spans the Hecke orbit, which implies $T_{p}\mathbb{K}_{2}(i/b,s_{i})$ can be written as a linear combination of $\mathbb{K}_{2}(h/b,s_{h})$ functions.
Next, denote the $n$th Fourier coefficient of $\mathbb{K}_{2}(i/b,s_{i})$ and $T_{p}\mathbb{K}_{2}(i/b,s_{i})$ by $A_{n}$ and $B_{n}$ respectively. Then we have the formula for $T_{p}$ on the Fourier coefficients
$B_{n}=A_{np}+p^{2}\chi(p)A_{n/p},$
where we assume $A_{n/p}$ is zero if $n/p$ is not an integer. Then, $A_{np}$ is nonzero only if $n\equiv k\mod b$. To see this, note that for $\mathbb{K}_{2}(j/b,s_{j})$ to be defined on a congruence group, the only possibilities for $b$ are $b=3,4,8,12,24$ based on Lemma 2.1. In all of these cases, $(\mathbb{Z}/b\mathbb{Z})^{\times}$ is isomorphic to a finite product of $\mathbb{Z}/2\mathbb{Z}$ with itself, and so each element has order 2. Recall $A_{n}$ is nonzero only if $n\equiv j\mod b$. Then for $A_{np}$ to be nonzero, we need $np\equiv j\mod b$ where $p\equiv i$. By assumption $ij=k$ in $(\mathbb{Z}/b\mathbb{Z})^{\times}$, and since each element is its own inverse, this implies $j=ik$. Since $p\equiv i\mod b$, we have $np\equiv j\mod b$ if and only if $n\equiv k\mod b$.
We now consider $A_{n/p}$. Analogous to above, since $p\equiv i\mod b$ and $i$ is its own inverse, $1/p$ is equivalent to $i$ in $(\mathbb{Z}/b\mathbb{Z})^{\times}$. Thus, for $n/p$ to be congruent to $j\mod b$, we must have $n\equiv ij=k\mod b$ as well. To conclude, by the first paragraph, $T_{p}\mathbb{K}_{2}(j/b,s_{j})$ is a linear combination of $\mathbb{K}_{2}(h/b,s_{h})$ functions, and since $\mathbb{K}_{2}(k/b,s_{k})$ is the only member of its family with nonzero coefficients for $n\equiv k\mod b$, we have proved the lemma. ∎
We establish the notation
$K_{i}:=\mathbb{K}_{2}(i/b,s_{i}),$
since the Lemma shows the action of the Hecke operators on a Galois family depends only on the index $i\in(\mathbb{Z}/b\mathbb{Z})^{\times}$.
As a corollary of Lemma 3.2, we can formulate the actions of the Hecke operators on this family $S_{r_{m},s_{m}}$ as follows.
Corollary 3.3.
Assume $S_{r_{m},s_{m}}=\{K_{i}\}_{i\in(\mathbb{Z}/b\mathbb{Z})^{\times}}$ is a Galois family. For any primes $p$ and $\ell$ coprime to $N$,
$T_{\ell}T_{p}=T_{p}T_{\ell}=D(p,\ell)T_{[p\ell]},\quad D(p,\ell)\in\mathbb{Z},$
where $[p\ell]\equiv p\ell\mod b$ is defined so that $0<[p\ell]<b$. Therefore, the Hecke algebra on this set is determined by $T_{p}$, $p\in(\mathbb{Z}/b\mathbb{Z})^{\times}$. Moreover, for a prime $p\in(\mathbb{Z}/b\mathbb{Z})^{\times}$, after a suitable reordering of the basis $S_{r_{i},s_{i}}$, the action of $T_{p}$ is given by
$\begin{pmatrix}T_{p,1}&0&\dots&0\\ 0&T_{p,2}&\ddots&0\\ 0&\ddots&\ddots&0\\ 0&\dots&0&T_{p,\frac{\phi(b)}{2}}\end{pmatrix},\quad T_{p,i}=\begin{pmatrix}0&a_{i}\\ b_{i}&0\\ \end{pmatrix},\quad a_{i}b_{i}=D(p,p).$
[…]
Proof.
Note
$T_{p}T_{\ell}K_{j}=C(\ell,j)C(p,[j\ell])K_{[p\ell j]}$
by standard properties of the Hecke operators and Lemma 3.2. On the other hand,
$T_{[p\ell]}K_{j}=C([p\ell],j)K_{[[p\ell]j]},$
since every integer less than $b$ and coprime to $b$ happens to be a prime number in all of our cases. But $[[p\ell]j]=[p\ell j]$ in $(\mathbb{Z}/b\mathbb{Z})^{\times}$, and so we have
$T_{\ell}T_{p}K_{j}=\frac{C(\ell,j)C(p,[j\ell])}{C([p\ell],j)}T_{[p\ell]}K_{j}.$
This constant does not depend on $j$ because the Hecke algebra is commutative, and it is by definition $D(p,\ell)$ as stated in the Corollary. The remainder of the corollary follows immediately from Lemma 3.2.
Theorem 3.4.
Assume $\{\mathbb{K}_{2}(i/b,s_{i})\}_{i\in(\mathbb{Z}/b\mathbb{Z})^{\times}}$ is a Galois family in the Hecke orbit of the newform $f$. Then there exist algebraic numbers $\beta_{i}$ integral over a quadratic number field depending on $i$ so that
$\sum_{i\in(\mathbb{Z}/b\mathbb{Z})^{\times}}\beta_{i}\mathbb{K}_{2}(i/b,s_{i})$
is an eigenvector of $T_{p}$ for all $p$ coprime to 6. Moreover, any Hecke eigenform constructed as above can be written as $f\otimes\phi$, where $\phi$ is some trivial or quadratic character with conductor dividing $b$.
Proof.
By Corollary 3.3, the eigenvalues of $T_{p}$ are ±$\pm\sqrt{D(p,p)}$ for each $p$. The only possible combination of $\mathbb{K}_{2}(1/b,s_{1})$ and $\mathbb{K}_{2}(i/b,s_{i})$ that is an eigenvector of $T_{i}$ with eigenvalue ±$\pm\sqrt{D(p,p)}$ is
$\mathbb{K}_{2}(1/b,s_{1})\pm\sqrt{D(i,i)}\mathbb{K}_{2}(i/b,s_{i}).$
Because the matrices for $T_{i}$ are simultaneously diagnolizable by the spectral theorem, there is a basis for $S_{i/b,s_{i}}$ which are eigenvectors for all the $T_{i}$. By the discussion above, each such eigenvector must have the form
$\sum_{i\in(\mathbb{Z}/b\mathbb{Z})^{\times}}\pm\sqrt{D(i,i)}\mathbb{K}_{2}(i/b,s_{i}).$
The coefficients $\beta_{i}$ in the statement are thus $\sqrt{D(i,i)}$ up to a sign. However, we cannot choose the sign arbitrarily according to Corollary 3.3; we must have that if $ij=k$ in $\mathbb{Z}$, then $\beta_{i}\beta_{j}=\beta_{k}$. We arbitrarily choose signs for the $\beta_{i}$ so that this is true.
The only ambiguity in constructing our $\beta_{i}$ is the choice of sign, and so any other Hecke eigenform constructed using the family $\{\mathbb{K}_{2}(i/b,s_{i})\}$ must have the same Fourier coefficients up to a sign. For the coefficients to remain multiplicative, the only way we can adjust this sign is equivalent to tensoring the coefficients by a quadratic character. In our linear combination, we only have control over the coefficients modulo $b$, and so the resulting character has to be on $(\mathbb{Z}/b\mathbb{Z})^{\times}$, that is, the conductor divides $b$
Because of (8), integrating both sides produces an exact $L$-value of the Hecke eigenform.
Corollary 3.5.
Assume $\{\mathbb{K}_{2}(i/b,s_{i})\}_{i\in(\mathbb{Z}/b\mathbb{Z})^{\times}}$ is a Galois family. Then there are algebraic numbers integral over a quadratic number field depending on $i$, $\beta_{i}$, so that for some quadratic character $\phi$ with conductor dividing $b$,
$L(f\otimes\phi,1)=\sum_{i\in(\mathbb{Z}/b\mathbb{Z})^{\times}}\beta_{i}F(i/b,s_{i}).$
Using the functional equation for a weight 3 Hecke eigenform, we can also obtain the special $L$-value at 2 from this formula as well.
Remark 4.
For a given newform $f$ of weight 3, note if $f$ has Hecke eigenvalue field larger than $\mathbb{Q}$, then the choice of $L$-function depends on which (inner) twist we choose in Corollary 3.5. We establish our choice of twist in Table 3, and henceforth, when we write $L(f,1)$, we are referring to the twist given in that table unless noted otherwise.
3.1.2. An Example of Constructing Hecke Eigenforms
To illustrate the Theorem above, we use the family that includes the pair $(1/8,5/8)$, in class 7 of Table 1 below. The computation (10) shows the family is Galois, and the conjugates have the form $(a/8,a/8+1/2)$ for $a=1,3,5,7$.
Using the level and character provided in [1] and carefully studying the Fourier coefficients, we determine that all four of these are in the Hecke orbit of $f_{64.3.d.a}$, where we use the LMFDB label to denote the Hecke eigenforms, which agrees with Table 1. To find the linear combination of the conjugates, it suffices to determine the action of the Hecke operators $T_{p}\mathbb{K}_{2}(1/8,5/8)$ for $p=3,5,7$. A computation on the Fourier coefficients reveals the following action.
$\begin{array}[]{c||c|c|c|c}&f_{1}&f_{3}&f_{5}&f_{7}\\ \hline\cr T_{3}&12f_{3}&f_{1}&-4f_{7}&-3f_{5}\\ T_{5}&-48f_{5}&16f_{7}&f_{1}&-3f_{3}\\ T_{7}&-64f_{7}&16f_{5}&-4f_{3}&f_{1}\end{array}$
From the table, we see that $T_{3}^{2}=12$, $T_{5}^{2}=-48$, and $T_{3}T_{5}=-3T_{7}$. We conclude $\beta_{1}=2\sqrt{3}$, $\beta_{2}=4i\sqrt{3}$ and because of the multiplicative nature of the coefficients, $\beta_{3}=-\beta_{1}\beta_{2}/3=-8i$. We find that
$f_{1/8,5/8}=f_{64.3.d.a}=\mathbb{K}_{2}(1/8,5/8)+2\sqrt{3}\mathbb{K}_{2}(3/8,7/8)+4\sqrt{3}i\mathbb{K}_{2}(5/8,9/8)-8i\mathbb{K}_{2}(7/8,11/8).$
By Corollary 3.5, the above formula also immediately transforms into the special $L$-value at 1 if we replace $\mathbb{K}_{2}$ with $F$. That is,
$L(f_{64.3.d.a},1)=F(1/8,5/8)-2\sqrt{3}F(3/8,7/8)-4\sqrt{3}iF(5/8,9/8)-8iF(7/8,11/8).$
We use this method repeatedly in the rest of the paper, especially when working with $L$-values. For convenience, the correct linear combination of conjugates for all the Galois cases is listed in the Appendix.
3.2. Twisting and the Kummer Transformation
Recall from the introduction that we have the following twisting property between families of conjugates.
$\mathbb{K}_{2}(1/4,1)=q-2q^{5}-7q^{9}+14q^{13}+18q^{17}-32q^{21}-21q^{25}+14q^{29}+O(q^{30})$
$\mathbb{K}_{2}(1/4,3/4)=q+2q^{5}-7q^{9}-14q^{13}+18q^{17}+32q^{21}-21q^{25}-14q^{29}+O(q^{30}).$
We interpret this as a modular forms version of Theorem 1.1 above in terms of families of conjugates. The relationship is summarized the diagram below.
In the diagram, the dotted lines denote relations between modular forms (Theorem 3.6 below), the dashed lines are relations between differentials on a hypergeometric surface (Definition 2.2), and the solid lines are transcendence relations of hypergeometric series (Theorem 1.1). Our application to $L$-values of newforms for some special cases arises because of a combination of all three of these aspects. In this section, we focus on the twisting. This is explained by symmetries of the hypergeometric functions underlying $\mathbb{K}_{2}(r,s)$.
Theorem 3.6.
Let $h=h(r,s)=r-s+3/2$. Then if $20<h(r,s)<3/$ 2, for any positive integer $n$, the $n$th Fourier coefficients of $\mathbb{K}_{2}(r,s)(\tau)$ and $\mathbb{K}_{2}(r,h)(\tau)$ are the same up to a sign. Moreover, if $a$ is the numerator of $r$, the $n$th coefficients are equal if and only if $n\equiv a\mod N$, where $N$ is as in (6).
[…]
Proof.
Apply the Pfaff transformation (see [2]),
$\,{}_{2}F_{1}\left[\begin{matrix}1/2&1/2\vskip 3.0pt plus 1.0pt minus 1.0pt\\ &1\end{matrix}\;;\;x\right]=(1-x)^{-1/2}\,_{2}F_{1}\left[\begin{matrix}1/2&1/2\vskip 3.0pt plus 1.0pt minus 1.0pt\\ &1\end{matrix}\;;\;\frac{x}{x-1}\right],$
to
$\mathbb{K}_{2}(r,s)(\tau)d\tau=\lambda^{r}(1-\lambda)^{s-r-1}\,_{2}F_{1}\left[\begin{matrix}1/2&1/2\vskip 3.0pt plus 1.0pt minus 1.0pt\\ &1\end{matrix}\;;\;\lambda\right]\frac{d\lambda}{2\pi i\lambda}.$
This gives us
[…]
Making a change of variable, $t=\lambda/(\lambda-1)$, and using that $\lambda(\tau+1)=\lambda/(\lambda-1)$ (see e.g. [40]), we compute
$\mathbb{K}_{2}(r,s)(\tau)d\tau=$ $\left(\frac{t}{t-1}\right)^{r}\left(\frac{1}{t-1}\right)^{s-r-3/2}\,_{2}F_{1}\left[\begin{matrix}1/2&1/2\vskip 3.0pt plus 1.0pt minus 1.0pt\\ &1\end{matrix}\;;\;t\right]\frac{dt}{2\pi it(t-1)}$
$=(-1)^{r}(1-t)^{1/2-s}t^{r}\,_{2}F_{1}\left[\begin{matrix}1/2&1/2\vskip 3.0pt plus 1.0pt minus 1.0pt\\ &1\end{matrix}\;;\;t\right]\frac{dt}{2\pi it}$
$=(-1)^{r}\mathbb{K}_{2}(r,r-s+3/2)(\tau+1)d\tau,$
invoking that the linear fractional transformation $\lambda/(\lambda-1)$ is its own inverse repeatedly. Because of this, the $q$-expansions of $\mathbb{K}_{2}(r,s)(\tau)$ and $\mathbb{K}_{2}(r,h)(\tau)$ differ by a root of unity, and as the $\mathbb{K}_{2}(r,s)$ functions always have rational coefficients, this means the Fourier coefficients must differ by a sign.
Let $r=a/b$ for $a$ and $b$ coprime integers. For the second part, the key observation is that $b$ is always equal to $N/2$, as noted after (6). Also, note that $(-1)^{r}=e^{-a\pi i/b}=e^{-2a\pi i/N}.$ We temporarily—only for the remainder of this proof—use the local uniformizer $q_{N}=e^{2\pi i\tau/N}$. The map $\tau\mapsto\tau+1$ sends
$e^{2\pi i\tau/N}\mapsto e^{2\pi i\tau/N}e^{2\pi i/N},$
and so $q_{N}^{n}\mapsto e^{2n\pi i/N}q_{N}^{n}$. Multiplying $e^{-2a\pi i/N}$ through, we see the $n$th Fourier coefficient for $\mathbb{K}_{2}(r,s)(\tau)$ is the same as the $n$th Fourier coefficient of $\mathbb{K}_{2}(r,h)(\tau)$ up to $e^{2(n-a)\pi i/N}$. Again writing out the $q_{N}$-expansion, we see the first term in the Fourier expansion of $\mathbb{K}_{2}(r,s)(\tau)$ is $q_{N}^{a}$, and all other non-zero coefficients are equivalent to $a$ modulo $b$. Since $b=N/2$, this implies $n=a+mN/2$ for some integer $m$. This divides the $n$ into two congruence classes mod $N$, equivalent to $a$ when $m$ is even and $a+N/2$ when $m$ is odd. Evidently, the signs do not switch if and only if $m$ is even, so this proves the claim. ∎
Note that this doesn’t even require our modular form to be defined on a congruence group, and the twist could be the same as the original modular form. However, the phenomenon is called twisting because in all but one of the Galois cases, the completed (as in Theorem 3.4) two $q$-series differ by a twist of a character. For example, the newform
(14) $f_{64.3.c.b}=\mathbb{K}_{2}(1/4,1)+4i\mathbb{K}_{2}(3/4,1)$
is related by Kummer to
$f_{32.3.c.a}=\mathbb{K}_{2}(1/4,3/4)+4i\mathbb{K}_{2}(3/4,5/4)$
and $f_{64.3.c.b}=\chi_{8.b}\otimes f_{128.3.c.a}$, where $\chi_{8.b}$ is also given by its LMFDB label. The one exception is Class 3 in Table 1, which occurs because one of the corresponding Hecke eigenforms is an oldform for the level specified in Lemma 2.1.
All Hecke stable families that are not Galois are self-twisted in the following sense.
Definition 3.7.
We say a family of conjugates $(r_{i},s_{i})$ is self-twisted if the Kummer twist of $(r_{i},s_{i})$ equals $(r_{j},s_{j})$ for some $j$ in the same family.
All the non-CM self-twisted families arise from applying the Atkin–Lehner involution $W_{2}$ to a Galois family. For example, the Atkin–Lehner involution maps $\mathbb{K}_{2}(a/8,a/8+1/2)$ to $\mathbb{K}_{2}(1/2,a/8+1/2)$, up to a constant. We can then easily check that the latter family is self-twisted. Each of these families has the same number of conjugates, so there should be a relation between them. This is discussed more in Section 4.1, and the specific relation for the family $(a/8,a/8+1/2)$ is Proposition 4.4. This allows us to compute the action of the Hecke operators indirectly.
3.3. Classification of Conjugate Families
In this section, we provide a classification of our families by enumerating all the possible families. Understanding our situation in terms of the group action of $D_{6}:=\langle A,K\rangle$ on $\mathbb{S}_{2}^{\prime}$ allows us to classify our modular forms in a natural way. We demonstrate this pictorially by constructing a graph where the nodes are the elements of $\mathbb{S}_{2}^{\prime}$. The edges are the pairs related by $A$ (colored red) or by $K$ (colored black), and a black dotted line if two vertices are in the same Hecke orbit. An example is given in Figure 1 below. For example, $(1/4,5/4)$ corresponds to a level 16 holomorphic modular form. However, applying $K$ to this produces $(3/4,3/2)$, which is not in $\mathbb{S}_{2}^{\prime}$. The figure shows first a specific connected component of the graph, and then a more general connected component.
Figure 1. The graph associated to class 6 on the left, with the Atkin–Lehner involution in red, the Kummer twist in black, and elements in the same Hecke orbit dotted. The graph on the right is associated to a generic non-CM four-term class
3.3.1. List of all Hecke Stable Classes
The main result of this section is Table 1 below. Each row of the table corresponds to a connected component of the graph described above, including all three types of edges. As a result of the table, we are able to find all connected components of the graph consisting of Hecke stable families.
Theorem 3.8.
There are exactly 29 Hecke stable families, falling into 11 connected components in the graph of Hecke stable families. Moreover, there are 12 non-CM Galois families falling into 6 connected components.
[…]
Definition 3.9.
For a representative pair $(r,s)$ in a given family of conjugates, we denote its stabilizer subgroup in $D_{6}$ by $\text{Stab}(r,s)$.
Note this is stabilized within the family. For example, we would consider $A$ to be in $\text{Stab}(1/8,1)$ because $A(1/8,1)=(7/8,1)$, which is conjugate to $(1/8,1)$. We can classify the stabilizer as follows.
Proposition 3.10.
For every pair $(r,s)$ in a Hecke stable family outside of Class 1 and 5, we have
$\text{Stab}(r,s)\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}.$
If $(r,s)$ is in Class 1 and 5, then
$\text{Stab}(r,s)\cong D_{6}.$
Thus, in the Galois cases,
$|D_{6}/\text{Stab}(r,s)|=3,$
and so Table 1 is divided into three columns. For classes 1 and 5, the stabilizer is all of $D_{6}$, which implies there is only one family of conjugates. Therefore, we put a star ()(*)( ∗ ) on the second and third columns, denoting that this family is the same as the previous one. Namely, the three columns for these classes denote the action of the Kummer twist and the Atkin–Lehner involution, respectively.
Label Data LMFDB K-Data K-LMFDB AL-K-Data CM?
[…]
2 $1/4,5/4$ 16.3.c.a $1/4,1/2$ 64.3.c.a $1,5/4$ yes
3 $1/3,7/6$ 36.3.d.a $1/3,2/3$ $\begin{array}[]{c}\text{36.3.d.a}\\ \text{36.3.d.b}\end{array}$ $1/6,5/6$ yes
4 $1/8,9/8$ 32.3.d.a $1/8,1/2$ 256.3.c.e $1,9/8$ yes
[…]
6 $1/4,3/4$ 32.3.c.a $1/4,1$ 64.3.c.b $1/2,3/4$ no
7 $1/8,5/8$ 64.3.d.a $1/8,1$ 256.3.c.g $1/2,5/8$ no
8 $1/8,7/8$ 128.3.d.c $1/8,3/4$ $\begin{array}[]{c}\text{256.3.c.c}\\ \text{256.3.c.f}\end{array}$ $1/4,5/8$ no
9 $1/12,11/12$ $\begin{array}[]{c}\text{288.3.g.a}\\ \text{288.3.g.c}\end{array}$ $1/12,2/3$ $\begin{array}[]{c}\text{576.3.g.d}\\ \text{576.3.g.h}\end{array}$ $1/6,7/12$ no
10 $1/24,17/24$ 288.3.b.c $1/24,5/6$ 2304.3.g.y $2/3,17/24$ no
11 $1/24,23/24$ 1152.3.b.i $1/24,7/12$ $\begin{array}[]{c}\text{2304.3.g.p}\\ \text{2304.3.g.w}\end{array}$ $1/12,13/24$ no
Table 1. Classes of Hecke Stable Families.
The K-Data and K-LMFDB in Table 1 signify the Kummer twist of the initial pair (not the action of the matrix $K$ or the Kummer transformation (Theorem 1.1), and the corresponding LMFDB label. By Theorem 3.4, we can construct any twist by a quadratic character of our initial modular form if $b$ divides the conductor. When this is an inner twist, the LMFDB label is the same, but when it is not an inner twist, the label is different. For completeness, we list both labels in these cases.
A complete list of all the conjugates in each Hecke stable family and how to combine them into a Hecke eigenform is provided in the Appendix. All of the Galois families are Hecke stable because we can check that Lemma 3.2 holds. The first and second columns for classes 1-4 and 6-11 are Galois. The third column corresponds to the action of the Atkin–Lehner involution on the first column.
Lemma 3.11.
$\mathbb{K}_{2}(r_{i},s_{i})$ is Hecke stable if and only if $\mathbb{K}_{2}(s_{i}-r_{i},s_{i})$ is as well.
Proof.
Because the Atkin–Lehner involution commutes with the Hecke operators, the Atkin–Lehner involution maps forms in the same Hecke orbit to forms in the same Hecke orbit, that is, if the family of $\mathbb{K}_{2}(r,s)$ is Hecke stable, then the family of $\mathbb{K}_{2}(s-r,s)$ is as well. ∎
Therefore, the third column for Classes 1-4 and 6-11 are Hecke stable. Class 5 is a special case. We will show separately that there are simple relations among the three pairs in Class 5 in [27], which proves the family is Hecke stable. To prove that Table 1 encompasses all possible cases, one can simply enumerate every other family and show they are not Hecke stable.
3.3.2. List of non-Hecke stable classes
Note that all the underlying Hecke eigenforms here are CM. Therefore, these cases will be addressed primarily in [27]. However, we write out the families to verify the classification in this paper.
Data LMFDB 1 LMFDB 2 K-Data K-LMFDB 1 K-LMFDB 2
12 $1/24,5/24$ 576.3.b.b 576.3.b.c $1/24,1/3$ 2304.3.g.t 2304.3.g.s
13 $1/12,17/12$ 36.3.d.b 144.3 g.c $1/12,1/6$ 576.3.g.c 576.3.g.f
14 $1/8,11/8$ 128.3.d.a 128.3.d.b $1/8,1/4$ 256.3.c.b 256.3.c.d
15 $1/24,35/24$ 1152.3.b.b 1152.3.b.f $1/24,1/12$ 2304.3.g.e 2304.3.g.i
Table 2. Classes that are not Hecke stable
As in the Hecke stable case, the right-hand side of the above table is mapped by the Atkin–Lehner involution to itself, and the the left hand side is mapped to a self-twisted family. We omit the details for the non-Hecke stable self-twisted families. However, if we count all of these pairs as well as those in our two tables, we determine the number of non-degenerate pairs is exactly 167, this size of $\mathbb{S}_{1}$, showing our classification is complete.
4. Part 2: Identities of L-values and Transcendence
In the first part of the paper, we have focused on the properties of $\mathbb{K}_{2}(r,s)$ as a modular form, demonstrating the action of the Hecke operators, finding a twisting result, and classifying the families. In the second half, we will primarily be concerned with the periods, $\pi\cdot F(r,s)$. In the process, we apply the results of Part 1 to find new proofs for a variety of three-term identities, which also imply our main transcendence result.
According to Beyer et al. Section 2.2 [8], three-term identities arise from the cosets of a certain normal subgroup of $S_{6}\times C_{2}$. In the Galois cases, one important subgroup is the commutator $D_{6}^{\prime}=[D_{6},D_{6}]$ in $D_{6}$. We have
$D_{6}/D_{6}^{\prime}\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}=\langle M_{1},M_{2}\rangle,$
where
(15) $M_{1}:=AKA,\quad M_{2}:=KAK.$
Using the matrix form of $A$ and $K$, it is straightforward to verify that these are coset representatives and $M_{2},$ $M_{1}$, and $M_{1}M_{2}$ are indeed involutions. Note that the Kummer twist fixes (within the family) the right-hand side of Table 1, and the Atkin–Lehner involution fixes the left-hand column of Table 1. This implies $M_{1}$ and $M_{2}$ fix the conjugates of the left-hand column of Table 1. Therefore, these operators correspond to switching to a different conjugate pair in these cases, analogous to the action of the Hecke operators on the forms $\mathbb{K}_{2}(r,s)$ in Lemma 3.2. Then Theorem 1.4 can be summarized for half of the cases as follows.
Theorem 4.1.
Let $\mathfrak{a}$ be the hypergeometric multiset $HD(r,s)$ associated to the left-hand column of families 4, 5 and 9 with minimal or maximal value of $r$. Then there are algebraic constants $\alpha_{1},\alpha_{2}$ so that
(16) $F(\mathfrak{a})=\alpha_{1}\cdot F(M_{1}\mathfrak{a})+\alpha_{2}\cdot F(M_{2}\mathfrak{a}).$
From a transcendence point of view, it is especially notable that the constants are algebraic – this is not guaranteed for any three-term identity.
4.1. Three-Term Hypergeometric Identities
In this section, we prove our first three-term identities, some of which will play a role in the proof of Theorem 1.4.
To illustrate, consider the family
$\{\mathbb{K}_{2}(1/2,3/4),\mathbb{K}_{2}(1/2,5/4)\},$
which is self-twisted. This family is obtained by applying the Atkin–Lehner involution $W_{2}$ to
$\{\mathbb{K}_{2}(1/4,3/4),\mathbb{K}_{2}(3/4,5/4)\}.$
Since both should span the space of of the same Hecke eigenform, we expect there to be relations between them. Using the Fourier coefficients, we can check that
$f_{32.3.c.a}$ $=\frac{\mathbb{K}_{2}(1/2,3/4)(4\tau)+\mathbb{K}_{2}(1/2,5/4)(4\tau)}{2}+i\frac{\mathbb{K}_{2}(1/2,3/4)(4\tau)-\mathbb{K}_{2}(1/2,5/4)(4\tau)}{2}$
$=\mathbb{K}_{2}(1/4,3/4)(8\tau)+2i\mathbb{K}_{2}(3/4,5/4)(8\tau)$
Since the real and imaginary parts are the same, we obtain the following identity of eta products.
Proposition 4.2.
Let $\tau\in\mathcal{H}$. Then
$\mathbb{K}_{2}(1/4,3/4)(8\tau)=\frac{\eta(8\tau)^{12}}{\eta(4\tau)^{2}\eta(16\tau)^{4}}$ $=\frac{1}{2}\left(\eta(2\tau)^{4}\eta(8\tau)^{2}+\frac{\eta(4\tau)^{12}}{\eta(2\tau)^{2}\eta(8\tau)^{4}}\right)$
$=\frac{\mathbb{K}_{2}(1/2,3/4)(4\tau)+\mathbb{K}_{2}(1/2,5/4)(4\tau)}{2}.$
Integrating both sides, we find the three-term identity
$F(1/4,3/4)/2={F(1/2,3/4)+F(1/2,5/4)}.$
The more general statement in terms of Coxeter groups is the following.
Theorem 4.3.
For $r=a/b$, $b=4,8,12,24$ and $a$ any integer coprime to $b$ and less than $b/2$, we have
$F(r,1-r)/2=F(AM_{1}(r,1-r))+F(AM_{2}M_{1}(r,1-r)).$
Note that $AM_{1}(r,1-r)=(2r,r+1/2)$ and $AM_{2}M_{1}(r,1-r)=(2r,r+1).$
Proof.
Set $s=1-r$. We will use that if $r=a/b$ where $a$ and $b$ are in lowest terms, then $N=2b$. First, note
$A(r,s)=(s-r,r)=(1-2r,r).$
Also, Theorem 1.1 maps
$K(1-2r,r)=(2r,r+1/2)$
. The Fourier expansions of $\mathbb{K}_{2}(2r,r+1/2)(b\tau)$ and $\mathbb{K}_{2}(r,1-r)(2b\tau)$ both have initial term $q$, and they are in the same Hecke orbit. We can then check in each case that $\mathbb{K}_{2}(2r,r+1/2)(b\tau)$ and $\mathbb{K}_{2}(2r,r+1)(b\tau)$ are Kummer twists of each other, and their sum is equal to $\mathbb{K}_{2}(r,1-r)(2b\tau)$. Also,
[…]
If we change the variable of $\mathbb{K}_{2}(2r,r+1/2)(b\tau)$ to $\mathbb{K}_{2}(2r,r+1/2)(2b\tau)$, that multiplies the integral by $2$. Multiplying through by $-2\pi i$ and integrating both sides of the above equality gives
[…]
$=-2\pi i\int_{0}^{i\infty}\frac{\mathbb{K}_{2}(2r,r+1/2)+\mathbb{K}_{2}(2r,r+1)}{2}(2b\tau)d(2\tau)=2(F(2r,r+1/2)+F(2r,r+1)).$
Numerical evidence suggests Theorem 4.3 holds for other positive integers $b$ in addition to the ones listed, but a new proof would be needed.
Remark 5.
There are also versions of these identities for pairs $(r^{\prime},s^{\prime})$ conjugate to $(1-r,r)$ where $r^{\prime}>1/2$ found using the two bases for the corresponding Hecke eigenform as above. We call these sister identities. They have a less uniform shape than the identity above, and so we will only state them when we need them.
There are always identities related to the Atkin-Lehner involution, though they sometimes have more than three terms. For instance, there is the following five-term identity associated to class 7.
Proposition 4.4.
$2F(1/8,5/8)=F(1/2,5/8)+F(1/2,11/8)+3F(1/2,7/8)+3F(1/2,9/8).$
Proof.
Similar to above, compute the identity using the $q$-expansion of the underlying $\mathbb{K}_{2}(r,s)(8\tau)$ functions and then integrate. ∎
4.2. From Kummer to Shimura
In this section, we will give the proof of the extension of Kummer’s theorem to $L$-values, stated precisely as Theorem 4.6 below, as well as Theorem 1.4. The following theorem is fundamental.
Theorem 4.5 (Shimura [30]).
Suppose $f$ is a newform of weight $k>2$, $0<m<k$ is an integer, and $\chi$ a finite order character so that $\chi(-1)^{m-1}=1$. Then $L(f,1)=A\cdot(2\pi i)^{m-1}g(\chi)^{-1}L(f\otimes\chi,m)$, where $g(\chi)$ is a Gauss sum and $\alpha=\alpha(f,m,\chi)$ is an algebraic number depending on $f,m,\chi$. Moreover, if $K_{f}$ is the field generated by the Hecke eigenvalues of $f$, then $\alpha$ lies in $K_{f}$, and if $\sigma$ is a nontrivial automorphism of $K_{f}$, then
$\alpha^{\sigma}=L(f^{\sigma},1)/[(2\pi i)^{m-1}g(\chi^{\sigma})^{-1}L(f^{\sigma}\otimes\chi^{\sigma},m)].$
We rely on Theorem 4.5 where $\chi$ is a quadratic character to prove Theorem 1.4. Remarkably, the identities resulting from Theorem 1.4 then allow us to explicitly write down the algebraic constant guaranteed by Theorem 4.5 for the non-quadratic twists that arise from the Kummer twist, Theorem 3.6. These formulas, like (3) from the introduction, interrelate the Kummer transformation, Theorem 3.6, and the Shimura’s Theorem 4.5 in an attractive way.
Once we have proved this, we find $L$-value relations arising from the Kummer twist Theorem 3.6 in the non-CM setting by using the Kummer Transformation 1.1, which states
$F(r,s)=2^{4-8r}\alpha_{r,s}F(1-r,1/2+q-r)$
in terms of $F(r,s)$, where
(17) $\alpha_{r,s}=\frac{\Gamma(r)\Gamma(s-r)^{2}}{\Gamma(1-r)\Gamma(s-1/2)^{2}}.$
In most cases, $\alpha_{r,s}$ does not appear to be algebraic. Among the congruence examples, the only times it is algebraic can be summarized by the fact
$F(r,1)=2^{-4+8r}\csc(\pi r)F(1-r,3/2-r)$
and
$F(r,1-r)=2^{-4r}\csc(\pi r)F(1-r,3/2-2r),$
for any rational $r$. In terms of transcendence, the algebraic condition means that the families, e.g., $F(r,1)$ and $F(r,r+1/2)$ give two different transcendence bases over $\bar{\mathbb{Q}}$ for the space of periods.
There are only five classes in Table 1 where $\alpha_{r,s}$ is algebraic, namely 6, 7, 8, 9, and 10, excluding degenerate cases where one of the beta factors is 0. Notably, these 5 classes are all non-CM.
The simplest non-CM example is class 6, which is already addressed in [1]. Using Corollary 3.5, we have
$L(f_{32.3.c.a},1)=F(1/4,3/4)+4iF(3/4,5/4)$
and
$L(f_{64.3.c.b},1)=F(1/4,1)-4iF(3/4,1).$
In this case, Theorem 1.1 takes the form
(18) $\begin{split}F(1/4,3/4)&=4\sin(\pi/4)F(3/4,1)\\ F(3/4,5/4)&=\frac{1}{4}\sin(3\pi/4)F(1/4,1).\end{split}$
The critical fact is then that $\sin(\pi/4)=\sin(3\pi/4)$ and so we can factor out $i\sin(\pi/4)$, proving that
(19) $L(f_{32.3.c.a},1)=-\sin(\pi/4)[-4iF(3/4,1)+F(1/4,1)]=\frac{\sqrt{2}i}{2}L(f_{64.3.c.b},1).$
[…]
$L(f_{32.3.c.a},1)=i\sin(\pi/4)\cdot L(f_{64.3.c.b},1).$
However, when the denominator of $r$ is e.g. 8 (classes 7 and 8), $\sin(\pi/8)$ does not equal $\sin(3\pi/8)$, and so we cannot use this method directly.
To generalize this to classes 7, 8, and 9, we use the (conjectural) transcendence basis over $\bar{\mathbb{Q}}$ from Theorem 1.4 and express the remaining two periods in terms of it. These families have extra structure from the underlying potentially 2-isotypic Galois representations, which explains the existence of the extra relations. After doing this, the Kummer transformation almost magically extends to an identity of $L$-functions as predicted by Shimura. This occurs because of the deep motivic meaning of the modularity result in [1] relating the two hypergeometric series to the two newforms, translated from the étale setting into the complex (de Rham) setting.
Theorem 4.6.
For the embeddings of the Fourier coefficients specified in Table 3, we have the relations
$L(f_{64.3.d.a},1)=-\frac{1}{2}\zeta_{48}\cdot L(f_{256.3.c.g},1)\quad|\quad L(f_{576.3.g.d},1)=\sqrt{3}\zeta_{8}^{7}\cdot L(f_{288.3.d.a},1)$
$L(f_{256.3.c.f},1)=\sqrt{2-\sqrt{2}-4i\sqrt{3-2\sqrt{2}}}\cdot L(f_{128.3.d.c},1)$
4.2.1. The Key Three-Term relations
We write the three-term identities in terms of four series for conciseness, but by taking the real and imaginary part of each side produces two three-term identities for each identity listed below. This proves Theorem 1.4.
Here is the first relation for class 7:
Lemma 4.7.
$F(1/8,5/8)-8iF(7/8,11/8)=(1-\sqrt{-2})(2F(3/8,7/8)+4iF(5/8,9/8)).$
Proof.
The newform $f_{64.3.d.a}$ has one even inner twist $\chi_{8.b}$, given by the quadratic Dirichlet character 8.b in the LMFDB. For this proof, take $f=f_{64.3.d.a}$. Then, since the twist is even, we have by Shimura’s result that $L(f\otimes\chi_{8.b},1)$ and $L(f,1)$ differ by an algebraic number. The Gauss sum of the Dirichlet character is $\sqrt{8}$, and so Shimura further predicts that $\sqrt{8}\alpha$ belongs to the Hecke eigenvalue field of $f$, which is $\mathbb{Q}(\zeta_{12})$. We may write $L(f\otimes\chi_{8.b},1)$ and $L(f,1)$ in terms of hypergeometric series using Corollary 3.5, as $\chi_{8.b}$ is a quadratic character with conductor dividing 16. Then, rearranging the relation $L(f,1)=\alpha L(f\otimes\chi_{8.b},1)$, we deduce
$F(1/8,5/8)-8iF(7/8,11/8)=\frac{\sqrt{3}(1-\alpha)}{1+\alpha}(2F(3/8,7/8)+4iF(5/8,9/8)).$
Now, let $\sigma$ denote the automorphism of $\mathbb{C}$ obtained by lifting the unique nontrivial automorphism of $\mathbb{Q}(\sqrt{3})$. By construction, $f^{\sigma}=f\otimes\chi_{8.b}$. The second part of Shimura’s theorem implies
$(\sqrt{8}L(f,1)/L(f\otimes\chi_{8.b},1))^{\sigma}=(\sqrt{8}\alpha)^{\sigma}=\sqrt{8}L(f^{\sigma},1)/L(f,1)=8/(\sqrt{8}\alpha).$
Let $\alpha^{\prime}:=\sqrt{8}\alpha$, an element of $\mathbb{Q}(\zeta_{12})$ by Shimura. In general, we may write elements of $\mathbb{Q}(\zeta_{12})$ as $\alpha^{\prime}=a+b\sqrt{3}$ for $a,b\in\mathbb{Q}(i)$, and then
$1/\alpha^{\prime}=\frac{a-b\sqrt{3}}{a^{2}-3b^{2}}.$
However, from above $\sigma(\alpha^{\prime})=a-b\sqrt{3}=8/\alpha^{\prime}$, and so $a^{2}-3b^{2}=8$. This implies the norm $N_{\mathbb{Q}(\zeta_{12})/\mathbb{Q}(i)}(\alpha^{\prime})=8$, and so the absolute norm is 64. Up to multiplication by a unit in the ring of integers, the only element of norm 64 in $\mathbb{Q}(\zeta_{12})$ is $2+2i$. This is because 2 is a divisor of 64, so 2 is ramified in $\mathbb{Q}(\zeta_{12})$ and factors as $(2)=(1+i)^{2}$. Moreover, the unit group of $\mathbb{Q}(\zeta_{12})$ has rank 1 and is generated by $\zeta_{12}-1$. Combining these two facts, we get that $[\alpha^{\prime}/(2+2i)]^{12}=(\zeta_{12}-1)^{n}$ for some integer $n$. The elements of the unit group lie in a lattice in Minkowski space, so we should be able to determine them numerically. In particular, let $B=|\zeta_{12}-1|$, where the bars denote complex modulus. Then we have
$\log_{B}|[\alpha^{\prime}/(2+2i)]^{12}|=n.$
Since $n$ is an integer and the integers are discrete in the reals, there is an open ball of radius, say, 1/4, around $n$ so that $n$ is the only integer in this interval. We can compute $\log_{B}|[\alpha^{\prime}/(2+2i)]^{12}|$ numerically to a high enough level of precision that we can determine it is within the ball of radius 1/4 around 12, and so we conclude $n=12$. Therefore, we have $\alpha^{\prime}/(2+2i)=(\zeta_{12}-1)\varepsilon$, where $\varepsilon$ is a 12th root of unity. The 12th roots of unity can be distinguished numerically, and so we check that $\varepsilon=-\zeta_{12}^{10}$. In conclusion, $\alpha^{\prime}=-(2+2i)(\zeta_{12}-1)\zeta_{12}^{10}=-2i(-1+\sqrt{3})$. From this, we can unravel the desired identity. ∎
There is a similar identity for the Kummer twist of the family above,
Lemma 4.8.
$F(1/8,1)+8iF(7/8,1)=-2(1-\sqrt{-2})(1+\sqrt{2})(F(3/8,1)-2iF(5/8,1)).$
Proof.
Apply the Kummer transformation to each side of the above Lemma, which in this case has the form $F(1-r,3/2-r)=\sin(\pi r)2^{4-8r}F(r,1),$ so we get
$\sin(\pi/8)(8F(7/8,1)+iF(1/8,1))=2\cos(\pi/8)(1-\sqrt{-2})(2F(5/8,1)+iF(3/8,1)).$
Multiply though by $i$ and note that $\cot(\pi/8)=1+\sqrt{2}$ and we arrive at the desired identity. ∎
The proof in class 8 is very similar to class 7.
Lemma 4.9.
$F(1/8,7/8)+8i\sqrt{2}F(7/8,9/8)=\zeta_{8}(2\sqrt{2}F(3/8,5/8)+4iF(5/8,11/8))$
Proof.
Similar to before, we twist $f_{128.3.d.c}$ by the character $\chi_{8.b}$, which is an even twist, so will give us an algebraic relation between the $L$-values. We have $g(\chi_{8.b})=\sqrt{8}$, and again let $\alpha$ be the constant relating the two $L$-values, and $\alpha^{\prime}=\sqrt{8}\alpha$. Take $\sigma$ to be the automorphism of $\mathbb{Q}(\zeta_{8})=\mathbb{Q}(\sqrt{2},i)$ which sends $\sqrt{2}\mapsto-\sqrt{2}$, $i\mapsto-i$. Then $f^{\sigma}=f\otimes\chi_{8.b}$ as above. From this, we can deduce that $\sigma(\alpha^{\prime})=8/\alpha^{\prime}$, and by the same argument, it follows that $\alpha^{\prime}$ has norm 64. We can take $2+2i$ again as the generator of the unique ideal of norm 64 in $\mathbb{Q}(\zeta_{8})$, since again, 2 is ramified. A fundamental unit is $u:=\zeta_{8}^{2}+\zeta_{8}+1$, and take $B=|u|$. Then we can compute $\log_{B}(|\alpha^{\prime}/(2+2i)|^{8})=8$ and using another numerical calculation, determine that $\alpha^{\prime}/(2+2i)=-u$. Simplifying, we get $\alpha^{\prime}=-2i(2+\sqrt{2})$, and then by rearranging the resulting equality of $L$-values we arrive at the identity above.∎
Lemma 4.10.
$F(1/8,3/4))-8\sqrt{-2}F(7/8,5/4)=\zeta_{8}^{3}(1-\sqrt{2})(4F(5/8,11/8)+2\sqrt{-2}F(3/8,5/8)).$
Proof.
The proof is identical to the proof of Lemma 4.8. ∎
The proof of the key lemma in class 8 requires several more steps, because there is no even inner twist of $f_{288.3.g.a}$. We still have a twist switching the signs of $5$ and $7$, but it is not an inner twist. Crucially, this means that we cannot use the second part of Shimura’s result, so the proof will be slightly different. We use on a result of [32], also given in [5] and [8] to handle this. We also rely on the following mild transcendence conjecture.
Conjecture 4.11.
$F(5/12,7/12)/F(7/12,17/12)$ is not an element of $\mathbb{Q}(\sqrt{3})$.
As discussed below, we expect that these two should even be algebraically independent, so this is not too surprising of a claim, but attempting to prove it is outside of the scope of this paper.
Proposition 4.12.
For any rational $r$ and $s$, we have that
$F(r,s)=-2^{6-4s-4r}\frac{\cos(\pi s)}{\cos\pi(r-1/2)}F(3/2-s,3/2-r)-2^{4s-8r+1}\frac{\sin(\pi r)}{\sin\pi(r-1/2)}F(s-r,s).$
Proof.
The identity is derived from a special case of a result of [5]: for a normalized ${}_{3}F_{2}$ we will denote $\hat{F}(a,b,c,d,e)$ as defined in Beyer et al. [8],
$\hat{F}(r,s)=\alpha_{1}\hat{F}(3/2-s,3/2-r)+\alpha_{2}\hat{F}({{r},{1+r-s},{r},{1/2+r},{1/2+r},})$
for explicit constants $\alpha_{1}$ and $\alpha_{2}$. Then we use a different result of Thomae, (see [2] Corollary 3.3.6), which for the normalized $\hat{F}$ directly tells us
$\hat{F}({{r},{1+r-s},{r},{1/2+r},{1/2+r}})=\hat{F}({{1/2},{1/2},{s-r},{1},{s}}).$
Rewriting this in terms of $F(r,s)$, we get to the identity. ∎
The identity becomes much simpler when $s=1-r$.
Corollary 4.13.
If $b$ is the denominator of $r$ and $r<b/2$, we have
$F(r,1-r)=-4F(r+1/2,3/2-r)+2^{5-12r}\tan(\pi r)F(1-2r,1-r).$
With Corollary 4.13, we are able to prove the key lemma of this section.
Lemma 4.14.
$F(1/12,11/12)+16iF(11/12,13/12)=\zeta_{3}^{2}(-4F(5/12,7/12)-4iF(7/12,17/12))$
Proof.
The twist switching $5$ and $7$ in this case is $\chi_{12.b}$ in the LMFDB, and we have $f_{288.3.g.a}\otimes\chi_{12.b}=f_{288.3.g.c}$. As before, if $L(f_{288.3.d.a},1)=\alpha L(f_{288.3.d.c},1)$, we can conclude
$F(1/12,11/12)+16iF(11/12,13/12)=\frac{1-\alpha}{1+\alpha}(-4F(5/12,7/12)-4iF(7/12,17/12)).$
We do know that $\sqrt{12}\alpha$ is in $\mathbb{Q}(i)$, which means $\alpha\in\mathbb{Q}(\sqrt{-3})=\mathbb{Q}(\zeta_{3})$.
Since applying the Atkin–Lehner involution to $F(1-2r,1-r)$ gives $F(r,1-r)$, we have already shown three term identities in our cases in Theorem 4.3. This means for any specific $r$, this can be rewritten as a 4-term identity. In our case, $r=1/12$, from Theorem 4.3 we have
$F(5/12,7/12)/2=F(5/6,11/12)+F(5/6,17/12)$
as well as its sister identity
$2F(11/12,13/12)=F(5/6,11/12)-F(5/6,17/12).$
Adding these up we get that
$2F(5/6,11/12)=F(5/12,7/12)+4F(11/12,13/12).$
Combined with Corollary 4.13, this produces the 4-term identity,
(20) $F(1/12,11/12)=-4F(7/12,17/12)+8\tan(\pi/12)[F(5/12,7/12)+4F(11/12,13/12)].$
We know from above that
$F(1/12,11/12)+16iF(11/12,13/12)=(a+b\sqrt{-3})(-4F(5/12,7/12)-4iF(7/12,17/12))$
for some rational $a$ and $b$. Taking the real and imaginary part of both sides, we obtain that
$F(1/12,11/12)=4aF(5/12,7/12)-4b\sqrt{3}F(7/12,17/12)$
and
$16F(11/12,13/12)=4aF(7/12,17/12)+4b\sqrt{3}F(5/12,7/12).$
If we put these into the identity (20), we get an equation only involving $F(5/12,7/12)$ and $F(7/12,17/12)$,
$-4aF(5/12,7/12)+4b\sqrt{3}F(7/12,17/12)=-4F(7/12,17/12)+$
$4\tan(\pi/12)[F(5/12,7/12)-16\sqrt{3}aF(7/12,17/12)-16bF(5/12,7/12)].$
This simplifies to
$4(-2+\sqrt{3}-a+(-3+2\sqrt{3})b)F(5/12,7/12)=4(1-(-2+\sqrt{3})a+\sqrt{3}b)F(7/12,17/12).$
If both of these constants are nonzero, that implies $F(5/12,7/12)=CF(7/12,17/12)$ for some constant $C$ contained in $\mathbb{Q}(\sqrt{3})$. Assuming Conjecture 4.11, this cannot be true, so both of these constants must be zero. This gives us a system of two equations in two variables, which by linear algebra has a unique solution, specifically $a=b=-1/2$. Since $-1/2-\sqrt{-3}/2=\zeta_{3}^{2}$, we are done. ∎
At least, the result for the Kummer twist is similar before.
Lemma 4.15.
$F(1/12,2/3)-16F(11/12,4/3)=1/4((-4+6i)+(2-4i)\sqrt{3})[-4F(5/12,4/3)+4iF(7/12,2/3)]$
Proof.
This is identical to the proof of Lemma 4.8. ∎
4.2.2. Extending Kummer’s Formula
Assuming the Lemmas of the previous section, we can now prove Theorem 4.6 via a straightforward by intricate computation. We illustrate with class 7. We omit the details for the other two cases, but the proof is extremely similar.
Theorem 4.16.
$L(f_{64.3.d.a},1)=-\frac{1}{2}\zeta_{48}\cdot L(f_{256.3.c.g},1).$
Proof of Theorem 4.16.
Using Lemma 4.7, compute
$L(f_{64.3.d.a},1)$
$=F(1/8,5/8)-2\sqrt{3}F(3/8,7/8)-4\sqrt{-3}F(5/8,9/8)-8iF(7/8,11/8)$
$=(F(1/8,5/8)-8iF(7/8,11/8))-2\sqrt{3}(F(3/8,7/8)+2iF(5/8,9/8))$
$=2(1-\sqrt{-2})(F(3/8,7/8)+2iF(5/8,9/8))-2\sqrt{3}(F(3/8,7/8)+2iF(5/8,9/8))$
$=2(1-\sqrt{-2}-\sqrt{3})(F(3/8,7/8)+2iF(5/8,9/8)).$
Then apply the Kummer transformation to get
$2(1-\sqrt{-2}-\sqrt{3})(F(3/8,7/8)+2iF(5/8,9/8))$
$=2(1-\sqrt{-2}-\sqrt{3})(2\cos(\pi/8)F(5/8,1)+i\cos(\pi/8)F(3/8,1))$
$=2i\cos(\pi/8)(1-\sqrt{-2}-\sqrt{3})(F(3/8,1)-2iF(5/8,1)).$
We now need to do the first half of this process in reverse using Lemma 4.8 to get the desired identity.
$L(f_{256.3.c.g},1)$
$=F(1/8,1)-2\sqrt{-3}F(3/8,1)-4\sqrt{3}F(5/8,1)+8iF(7/8,1)$
$=(F(1/8,1)+8iF(7/8,1))-2\sqrt{-3}(F(3/8,1)-2iF(5/8,1))$
$=-2(1-\sqrt{-2})(1+\sqrt{2})(F(3/8,1)-2iF(5/8,1))-2\sqrt{-3}(F(3/8,1)-2iF(5/8,1))$
$=-2((1-\sqrt{-2})(1+\sqrt{2})+\sqrt{-3})(F(3/8,1)-2iF(5/8,1)).$
Putting the two computations together, we get the relation
$L(f_{64.3.d.a},1)=\frac{2i\cos(\pi/8)(1-\sqrt{-2}-\sqrt{3})}{-2((1-\sqrt{-2})(1+\sqrt{2})+\sqrt{-3})}L(f_{256.3.c.g},1).$
Simplifying this constant and dividing to the other side of the equation gives us the desired result. ∎
4.3. Transcendence
For classical ${}_{2}F_{1}(z)$ functions, the transcendence of special values is fairly well understood. When the hypergeometric series can be written as an algebraic function of $z$, the transcendence is straightforward: ${}_{2}F_{1}(z)$ is algebraic if $z$ is algebraic and transcendental if $z$ is not. This is equivalent to the local system having finite monodromy, and these cases are classified in Beukers and Heckman [7], including for generalized hypergeometric series. The question of transcendence is harder for hypergeometric systems with infinite monodromy. In this case, it is well known that the monodromy group is isomorphic to a triangle group (see, e.g., [40]). The set of algebraic numbers $z$ for a fixed datum so that ${}_{2}F_{1}(z)$ is algebraic is called the exceptional set, denoted $\mathcal{E}$. Then Wolfart [35] and Cohen–Wüstholz [13] showed that when the monodromy group is an arithmetic triangle group, as classified by Takuechi [31], then the exceptional set is infinite; otherwise, it is finite. Wolfart’s proof uses the work of Wüstholz discussed in the introduction, and work of Cohen and Wüstholz relies on a special case of the André–Oort conjecture, which was proven by Edixhoven and Yafaev [15]. This construction associates each point element of the exceptional set to a particular abelian variety, which necessarily has CM type. However, the process is still rather theoretical, and so explicit elements of the exceptional set are nontrivial. For example, Yang in [38] proved an algebraic evaluation of a ${}_{2}F_{1}$ at a CM point using the action of the Hecke operators, and in [39] provides more examples using Borcherds lifts.
The proof for ${}_{2}F_{1}$ relies on two very deep theorems: Wüstholz’s theorem and a special case of the Andre-Oort conjecture. Moving up in dimension, we expect the geometry to become even more difficult, and so proving the transcendence of $F(r,s)$ without CM seems out of reach with our current knowledge. Note, we are interested in the transcendence degree of the periods for $H_{r,s}^{2,0}$. There are some results in the function field setting which indicate there might be similar transcendence results classically, e.g., [10], where the cohomology is nicer. However, transcendence results over $\bar{\mathbb{Q}}$ are generally very hard and require deep geometric theorems to prove. For example, $\Gamma(1/4)$ and $\Gamma(1/3)$ are known to be transcendental by Chudnovsky [12] using CM theory, but $\Gamma(1/5)$ is still unknown. A much stronger conjecture attributed to Rohrlich in [34] states that the only additive relations over ¯$\bar{\mathbb{Q}}(\pi)$ between
$\log\Gamma(j/5),\text{ }j=1,2,3,4$
arise from the well-known multiplicative formulas for the gamma function. In the same spirit, we conjecture the following.
Conjecture 4.17.
Let $f_{1},f_{2},f_{3},f_{4}$ be any of the four conjugate hypergeometric series used in the six Galois families from classes 7, 8, and 9 above. Then the only linear relations over $\bar{\mathbb{Q}}$ between these 4 can be derived from those given in Lemmas 4.7, 4.8, 4.9, 4.10, 4.14, and 4.15 respectively. Namely, the dimension of the vector space over $\bar{\mathbb{Q}}$ generated by $f_{1},f_{2},f_{3},f_{4}$ is 2.
Due to the discussion at the end of Appendix II, we can interpret this conjecture as a higher dimensional analogue of Wüstholz’s theorem in the special case that the endomorphism algebra has QM. The Hodge numbers of the underlying hypergeometric motive have the form $(\varphi(M),0,\varphi(M))$ over $\mathbb{Q}$. Then by a theorem of Totaro [33] the cohomology groups do not arise from the cohomology of Abelian variety, and so this conjecture does not follow from Wüstholz’s theorem. For example, if
$f_{1}=F(1/12,11/12),f_{2}=F(5/12,7/12),f_{3}=F(7/12,17/12),f_{4}=F(11/12,13/12),$
then taking the real and imaginary parts of Lemma 4.8 implies $f_{1}$ and $f_{4}$ can be written as a $\bar{\mathbb{Q}}$-linear combination of $f_{2}$ and $f_{3}$. Note that this implies Conjecture 4.11 above is true, though the current version is slightly stronger. However, as noted above, even proving that $f_{2}$ and $f_{3}$ are transcendental is likely to be very difficult, much less their linear independence. The best we can do is Theorem 1.4.
5. Further Directions
We covered all the non-CM classes except 10 and 11, which both have degree 8 Hecke eigenvalue fields, so are harder to work with. For class 11, the Kummer twist is odd, and so we cannot expect an algebraic relation between the $L$-values. However, we expect the main results of Part II to hold in these cases otherwise. There are also other hypergeometric data with which our methods could apply. For example, recently the author has found similar relations for certain ${}_{2}F_{1}(1)$ series in [28]. In addition, ${}_{2}F_{1}(z)$ hypergeometric functions are often modular forms on their underlying monodromy group, and then we can use the integral formula to obtain special $L$-values in the same manner as above. The datum $HD_{3}:=\{\{1/3,2/3\},\{1,1\}\}$ is the most approachable of these, since ${}_{2}F_{1}(HD_{3},z)$ is known to be a weight 1 modular form on $\Gamma_{0}(3)$, and has a classical expression as an eta quotient (see Borwein and Borwein [9]). The corresponding $\mathbb{K}_{3}$ function is mentioned in [1], and its further properties will be studied in a forthcoming paper of Grove [20]. More generally, Yang [37] has shown that modular forms on arithmetic triangle groups can be written as a combination of ${}_{2}F_{1}(z)$ functions. However, since in general there are no cusps and thus no $q$-expansions or eta products, these cases likely would be much harder.
6. Appendix I: The table of Hecke Eigenforms built from $\mathbb{K}_{2}$
In the appendix, we list all the completed modular forms we obtain using this method in Table 3. Note 6 of these are already given in [1], but we repeat them here for completeness. In cases where we are unable to express a full Hecke eigenform as a linear combination of the conjugates, we simply list the pairs. Observe also that changing the $\mathbb{K}_{2}(r,s)$ in this table to $F(r,s)$ will also give $N$ times the $L$-value of the eigenform, and so the table also gives us the special $L$-values obtained in the Hecke stable families, though in some cases these can be simplified. We omit the self-twisted families, as well as class 2. The latter, as well as class 1, only has a single conjugate for each family provided in the table. Class 5 will be addressed in detail in [27]. Finally, the remaining self-twisted families arise from the Atkin-Lehner involution applied to a family in Table 3. The conjugates may be obtained using Lemma 3.11, and constructing the eigenform involves simply using $q$-series identities like those in the proof of Theorem 4.3.
Label Eigenform
3.a $f_{36.3.d.a}=\mathbb{K}_{2}(1/3,7/6)$+2 $\mathbb{K}_{2}(2/3,5/6)$
3.b $f_{36.3.d.a}=\mathbb{K}_{2}(1/3,2/3)+2\mathbb{K}_{2}(2/3,4/3)$
4.a $f_{32.3.d.a}=\mathbb{K}_{2}(1/8,9/8)+2\mathbb{K}_{2}(3/8,11/8)$
4.b $f_{256.3.c.e}=\mathbb{K}_{2}(1/8,1/2)+2\mathbb{K}_{2}(3/8,1/2)$
6.a $f_{32.3.c.a}=\mathbb{K}_{2}(1/4,3/4)+4i\mathbb{K}_{2}(3/4,5/4)$
7.a $f_{64.3.d.a}=\mathbb{K}_{2}(1/8,5/8)-2\sqrt{3}\mathbb{K}_{2}(3/8,7/8)$
$-4\sqrt{-3}\mathbb{K}_{2}(5/8,9/8)-8i\mathbb{K}_{2}(7/8,11/8)$
7.b $f_{256.3.c.g}=\mathbb{K}_{2}(1/8,1)-2\sqrt{-3}\mathbb{K}_{2}(3/8,1])$
$-4\sqrt{3}\mathbb{K}_{2}(5/8,1)+8i\mathbb{K}_{2}(7/8,1)$
8.a $f_{128.3.d.c}=\mathbb{K}_{2}(1/8,7/8)+2\sqrt{2}\mathbb{K}_{2}(3/8,5/8)$
$+4i\mathbb{K}_{2}(5/8,11/8)+8\sqrt{-2}\mathbb{K}_{2}(7/8,9/8)$
8.b $f_{256.3.c.c}=\mathbb{K}_{2}(1/8,3/4)+2\sqrt{-2}\mathbb{K}_{2}(3/8,5/4)$
$+4\mathbb{K}_{2}(5/8,3/4)-8\sqrt{-2}\mathbb{K}_{2}(7/8,5/4)$
9.a $f_{288.3.g.a}=\mathbb{K}_{2}(1/12,11/12)-4\mathbb{K}_{2}(5/12,7/12)$
$-4i\mathbb{K}_{2}(7/12,17/12)+16i\mathbb{K}_{2}(11/12,13/12)$
9.b $f_{576.g.d}=\mathbb{K}_{2}(1/12,11/12)+4\mathbb{K}_{2}(5/12,7/12)$
$+4i\mathbb{K}_{2}(7/12,17/12)+16i\mathbb{K}_{2}(11/12,13/12)$
10.a $f_{1152.3.b.i}=-16\sqrt{-2}\mathbb{K}_{2}(23/24,25/24)+8\sqrt{14}\mathbb{K}_{2}(19/24,29/24)$
$+8\sqrt{7}\mathbb{K}_{2}(17/24,31/24)-4i\mathbb{K}_{2}(13/24,35/24)+4\sqrt{2}\mathbb{K}_{2}(11/24,13/24)$
$+2\sqrt{-14}\mathbb{K}_{2}(7/24,17/24)-2\sqrt{-7}\mathbb{K}_{2}(5/24,19/24)+\mathbb{K}_{2}(1/24,23/24)$
10.b $f_{2304.g.p}=16\sqrt{-2}\mathbb{K}_{2}(23/24,17/12)+8\sqrt{-14}\mathbb{K}_{2}(19/24,13/12)$
$+8\sqrt{7}\mathbb{K}_{2}(17/24,11/12)-4i\mathbb{K}_{2}(13/24,7/12)+4\sqrt{-2}\mathbb{K}_{2}(11/24,17/12)$
$+2\sqrt{-14}\mathbb{K}_{2}(7/24,13/12)-2\sqrt{7}\mathbb{K}_{2}(5/24,11/12)+\mathbb{K}_{2}(1/24,7/12)$
11.a $f_{288.3.b.c}=\mathbb{K}_{2}(1/24,17/24)-2\sqrt{6}\mathbb{K}_{2}(5/24,13/24)-2\sqrt{-15}\mathbb{K}_{2}(7/24,23/24)$
$-4\sqrt{10}\mathbb{K}_{2}(11/24,19/24)+4\sqrt{-15}\mathbb{K}_{2}(13/24,29/24)-8\sqrt{10}\mathbb{K}_{2}(17/24,25/24)$
$-8\mathbb{K}_{2}(19/24,35/24)-16\sqrt{-6}\mathbb{K}_{2}(23/24,31/24)$
11.b $f_{2304.3.g.y}=\mathbb{K}_{2}(1/24,5/6)-2\sqrt{6}\mathbb{K}_{2}(5/24,7/6)+2\sqrt{-15}\mathbb{K}_{2}(7/24,5/6)$
$+4\sqrt{-10}\mathbb{K}_{2}(11/24,7/6)+4\sqrt{15}\mathbb{K}_{2}(13/24,5/6)+8\sqrt{10}\mathbb{K}_{2}(17/24,7/6)$
$+8i\mathbb{K}_{2}(19/24,5/6)-16\sqrt{-6}\mathbb{K}_{2}(23/24,7/6)$
Table 3. Hecke Eigenforms constructed from Hecke stable families
Note for 3.a, we actually get an oldform of level 72, so the coefficients that are divisible by 2 are not consistent with the LMFDB, but after applying a Hecke operator $T_{p}$ with $p$ coprime to the level, they become the same. Also note that the formula for AL-3 only gives $\chi_{-4}f_{36.3.d.a}$, so some of the coefficients divisible by 2 are missing. Because this differs by an even twist, the $L$-value obtained from AL-3 and 3.a or 3.b differ by an algebraic number by Shimura’s theorem, which actually is an integer. Finally, for AL-11, there may be some way to combine these conjugates to get $f_{288.3.b.c}$, parallel to the difficult construction for AL-7, which is related to Proposition 4.4. However, this is tedious, and we leave it to the interested reader to find such a relation.
7. Appendix II: Geometric Considerations
Our motivation for the transcendence results in this paper has been focused on the two-dimensional Galois representations attached to the hypergeometric datum. This is convenient, because the Galois representations are well-understood from [1]. From a motivic point of view, we are using étale realization of the motive to motivate facts about the de Rham aspect. However, the disadvantage is that for general hypergeometric motives, the correspondence between behavior of the Galois representations and the behavior of the transcendence degree is less well-understood. Therefore, in this section, we outline how our transcendence result can be understood using purely complex geometry in a way that aligns better with the statement of Wüstholz’s theorem.
The modular form $\mathbb{K}_{2}(r,s)$ is not a differential itself. However, its period corresponds to a differential in a natural way. The base varieties for these differentials are so-called hypergeometric varieties. For nondegenerate cases, the varieties can be defined (see e.g. [26]) as
$y^{M}=Y_{r,s}:=x_{1}^{M/2}(1-x_{1})^{M/2}x_{2}^{M(1-r)}(1-x_{2})^{M(s-r)}(1-x_{1}x_{2})^{M/2},$
assuming that $r$ and $s-r$ are between 0 and 1. Then by the integral formula for hypergeometric series, we have
$\int_{0}^{1}\int_{0}^{1}\frac{dx_{1}dx_{2}}{y}=\int_{0}^{1}\int_{0}^{1}\frac{dx_{1}dx_{2}}{\sqrt[M]{Y{r,s}}}=P(r,s).$
The variety is badly singular for the locus $(0,x_{1},1/x_{1})$, i.e. when $x_{1}x_{2}=1$. Because these are surfaces over $\mathbb{C}$, there is a minimal resolution of the singularities, which we will denote by $X_{r,s}$. On a related note, an explicit partial compactification for these surfaces and higher dimensional analogues in the generic case $\lambda\neq 0,1$ is introduced by Kelly and Voight [22], which allows them to construct a Hodge structure on the surface. Moreover, the cohomology of the surface, based on their computation of the zeta function, arises either from hypergeometric motives or class field theory.
We will restrict our attention to a certain subspace of the de Rham cohomology, $H^{2}_{dR}(X_{r,s},\mathbb{C})$ corresponding to a hypergeometric motive. We note that there is and action of $\mu_{M}$, the $M$th roots of unity, on the space $H^{2,0}(X_{r,s})$ in Kelly and Voight’s paper, the compactification is equivariant with respect to this action. Throughout, we will use the notation $\zeta_{n}=e^{2\pi i/n}$. Parallel to the cases in e.g. [3], the action is given explicitly by $(x_{1},x_{2},y)\mapsto(x_{1},x_{2},\zeta_{M}^{-1}y)$ on $y^{M}=\alpha_{r,s}$. This induces an action on $H^{2}_{dR}(X_{r,s},\mathbb{C})$, and decomposes the cohomology into eigenspaces:
$H^{2}_{dR}(X_{r,s},\mathbb{C})\otimes_{\mathbb{Z}}\mathbb{Q}=\bigoplus_{n=0}^{M-1}V_{n}.$
By Hodge theory, the de Rham cohomology decomposes as
$H^{2}_{dR}(X_{r,s},\mathbb{C})\cong H^{2,0}(X_{r,s})\oplus H^{1,1}(X_{r,s})\oplus H^{0,2}(X_{r,s}),$
where $H^{p,q}(X_{r,s})$ denotes the Dolbeault cohomology. This induces a decomposition $V_{n}\cong V_{n}^{2,0}\oplus V_{n}^{1,1}\oplus V_{n}^{0,2}$.
Recall by Dolbeault’s theorem, $H^{2,0}(X_{r,s})$ is isomorphic to the space of holomorphic differential 2-forms. We construct differentials in $H^{2,0}(X_{r,s})$ following the procedure of Wolfart [35], Archinard [3] and others, for generalized Legendre curves. These differentials are constructed on $y^{M}=\alpha_{r,s}$ as
$\omega_{n,\mathfrak{a}}=\frac{x_{1}^{a_{1}}(1-x_{1})^{a_{2}}x_{2}^{a_{3}}(1-x_{2})^{a_{4}}(1-x_{1}x_{2})^{a_{5}}\cdot dx_{1}dx_{2}}{y^{n}},$
where $\mathfrak{a}=(a_{1},a_{2},a_{3},a_{4},a_{5})$. We then obtain a differential on $X_{r,s}$ by pulling back along the desingularization map - by an abuse of notation, we will use write $\omega_{n,\mathfrak{a}}$ for a differential on $X_{r,s}$ as well. The pullback of these differentials will not generally be holomorphic, except for certain choices of $\mathfrak{a}$. For generalized Legendre curves, Archinard [3] writes down the conditions on $\mathfrak{a}$ explicitly using the desingularization map. In our setting, the condition that $(r,s)\in\mathbb{S}_{1}$ will be a sufficient condition for the differential to be holomorphic.
To set this up, recall by Lemma 3.1, there exists a rational number $s_{1}$ so that $(1/b,s_{1})\in\mathbb{S}_{1}$ is conjugate to $(r,s)$. If there are multiple such numbers, we choose $s_{1}$ to be minimal. Therefore, there exists an integer $c$ coprime to $M$ and unique (up to the choice of $r,s$ and $c$) integers $k$ and $h$ so that $r=c\cdot 1/b+k$ and $s=c\cdot s_{1}+h$. By elementary number theory similar to the proof of Lemma 3.1, it is possible to choose a unique $c$ so that $-M<c<M$ assuming $r<1$.
Lemma 7.1.
If $(1/b,s_{1})$ and $(r,s)$ are as above and assuming that $r<1$, then the period of $\mathbb{K}_{2}(r,s)$ is equal the period $\omega_{c,\mathfrak{a}}$ up to $\pi$ times an algebraic number if $c\geq 0$ and is equal to the period of $\omega_{M+c,\mathfrak{a}}$ if $c<0$ otherwise. That is, conjugate $\mathbb{K}_{2}(r,s)$ correspond to periods defined on the same algebraic surface. Moreover, $\omega_{n,\mathfrak{a}}$ lies in the subspace $V_{n}$.
Proof.
Let $y=\sqrt[M]{C(1/b,s_{1})}$ and $y^{\prime}=\sqrt[M]{\alpha_{r,s}}$. By construction, $dx_{2}dx_{2}/y^{\prime}$ is equal to $\mathbb{K}_{2}(r,s)$ up to an algebraic constant. To show this is also a differential on $y^{M}=C(1/b,s_{1})$, we need to find a vector $\mathfrak{a}=(a_{1},a_{2},a_{3},a_{4},a_{5})$ so that
(21) $\frac{dx_{1}dx_{2}}{y^{\prime}}=\frac{x_{1}^{a_{1}}(1-x_{1})^{a_{2}}x_{2}^{a_{3}}(1-x_{2})^{a_{4}}(1-x_{1}x_{2})^{a_{5}}dx_{1}dx_{2}}{y^{c}}.$
By substituting $r=c\cdot 1/b+k$ and $s=c\cdot s_{1}+h$ and manipulating the differential, we find that the above is true for
$\mathfrak{a}=\left(\frac{c-1}{2},\frac{c-1}{2},ch-ck+(1/b-s_{1})(1-c^{2}),1+c-ck-\frac{1-c^{2}}{b},\frac{c-1}{2}\right).$
We know $c,h$ and $k$ are integers already. Since $M$ is even, $(c-1)/2$ is always an integer, as $c$ is coprime to $M$. Moreover, it is easy to check that $M\mid 1-c^{2}$ for all $c$ coprime to $M$ for $M=2,4,6,8,12,24$, the only possibilities for $M$ in our setting. Therefore, the entries of $\mathfrak{a}$ are integers. When $c\geq 0$, we are done. If $c\leq 0$, note that $y^{M}$ is a polynomial by definition. Multiplying the top and bottom of the right hand side in equation (21) and adjusting the $a_{i}$ using $y^{M}$ gives a differential of the form $\omega_{M+c,\mathfrak{a}}$ as desired.
The last claim is immediate from the definition of $\omega_{n,\mathfrak{a}}$, as clearly multiplying $y$ by $\zeta_{M}^{-1}$ will produce the eigenvalue $\zeta_{M}^{n}$. ∎
Remark 6.
The condition $M\mid 1-c^{2}$ does not hold for an arbitrary choice of $M$ and $c$. For instance, even for $M=16$, we could take $c=3$, but clearly 16 does not divide $83^{2}-1=$ 8. The underlying issue is that the definition of conjugates was essentially lifted from the ${}_{2}F_{1}(\lambda)$ setting, where Lemma 7.1 follows from a classical result about Riemann surfaces (refer to e.g. [3]). Mostly by luck, the same concept holds for our surfaces, but this does not work for general hypergeometric surfaces. To our knowledge, an explicit description of the complete space of holomorphic differentials and their periods on hypergeometric surfaces remains an open problem.
There are no examples where $r=a/b>1$ in the non-CM examples. However, these cases are also easily handled with a slight modification.
Theorem 7.2.
A non-CM family is Galois if and only if for all $n$ satisfying $(n,M)=1$, we have $V_{n}^{2,0}$ contains exactly one differential of the shape $\omega_{n,\mathfrak{a}}$. Moreover, for a Galois family, $\dim_{\mathbb{Q}}V_{n}^{2,0}=1$ and is generated by $\omega_{n,\mathfrak{a}}$.
Proof.
The converse is straightforward - the one holomorphic differential, $\omega_{i,\mathfrak{a}}$, in $V_{i}^{2,0}$, has period $P(i/b,s_{i})$ for some $s_{i}$ by the definition and the integral formula for hypergeometric series. Each of these corresponds to a modular form $\mathbb{K}_{2}(i/b,s_{i})$. If $\mathbb{K}_{2}(i^{\prime}/b,s_{i^{\prime}})$ is conjugate to $\mathbb{K}_{2}(i/b,s_{i})$, then it corresponds to some differential $\omega_{i^{\prime},\mathfrak{a}}$ by the discussion above. But there is only one such differential in each $V_{n}$, and so we must have $i\neq i^{\prime}$ for all conjugates. This implies the family is Galois.
Now assume $\{\mathbb{K}_{2}(i/b,s_{i})\}_{i\in(\mathbb{Z}/b\mathbb{Z})^{\times}}$ is a Galois family. Using Table 1, we can easily check that for all such families, $s_{i}>1/2$. Then by the Lemma above, this implies that each pair $(i/b,s_{i})$ corresponds to a differential $\omega_{i,\mathfrak{a}}$. But by the definition of Galois, each of the $i$ is distinct, and so there is exactly one $\omega_{i,\mathfrak{a}}$ in the space $V_{i}$. A result of Fedorov [16] allows us to compute the Hodge numbers for hypergeometric motives over $\mathbb{Q}(\zeta_{M})$ for $\lambda\neq 1$. In all of the Galois cases, these Hodge numbers are $(1,1,1)$, since $s_{i}>1/2$. When moving to $\lambda=1$, the rank is reduced by one in the middle cohomology, making the Hodge numbers $(1,0,1)$. Due to the correspondence above, we can assume that $\omega_{1,\mathfrak{a}}$ is in the cohomology associated to the hypergeometric motive, and so we deduce that $\dim_{\mathbb{Q}(\zeta_{M})}H_{r,s}^{2,0}=1$. Therefore, the dimension over $\mathbb{Q}$ is precisely $\varphi(M)$, corresponding to the $\varphi(M)$ holomorphic conjugates. There is exactly one conjugate in each $V_{n}^{2,0}$ by above, and there are precisely $\varphi(M)$ subspaces $V_{n}^{2,0}$. This forces $\dim V_{n}^{2,0}=1$. ∎
By the same reasoning as the proof above, we can conclude that $\dim_{\mathbb{Q}}H_{r,s}=2\varphi(M)$. The sub-Hodge structure $V_{n}\oplus V_{M-n}$ therefore has rank 4 over $\mathbb{Q}$ and Hodge numbers $(2,0,2)$. We conjecture the following.
Conjecture 7.3.
The endomorphism algebra $\text{End}_{hdg}(V_{n}\oplus V_{M-n})$ is isomorphic to a quaternion algebra over $\mathbb{Q}$.
This conjecture is likely true, since Totaro [33] has shown that Hodge structures of type $(2,0,2)$ admitting endomorphisms by an imaginary quadratic field satisfy the conclusion of the Conjecture. The only other possibility with extra endomorphisms, again due to Totaro, is a CM structure by a CM field of degree 4, which appears very unlikely numerically. We will leave the precise proof of the Conjecture to future work.
One could make a guess that any Hodge structure admitting endomorphisms by a quaternion algebra over $\mathbb{Q}$ has transcendence degree for periods of the first kind bounded by 2, parallel to Wüstholz’s theorem. We do not have sufficient evidence in this paper to state this as a proper conjecture, but there could be potential future work in this area as well. To our knowledge, the only existing literature in this area is related to Hodge structures with CM and the Gross-Deligne conjecture, see e.g. [4].